### v2

# Jacquet modules of Tate cohomology and base change lifting

Abstract.
Let $G$ be a connected reductive group defined over a non-Archimedean local field $F$ of residue characteristic $p$. Let $\ell$ be a prime number distinct from $p$. Let $E$ be a cyclic Galois extension of $F$ with $[E:F]=\ell$. Let $\Pi$ be a finite length $\overline{\mathbb{F}}_{\ell}$-representation (or an $\ell$-modular representation) of $G(E)\rtimes{\rm Gal}(E/F)$. In this context, we prove a conjecture of Treumann and Venkatesh which predicts that the Tate cohomology groups $\widehat{H}^{i}({\rm Gal}(E/F),\Pi)$ are finite length representations of $G(F)$. We discuss the explicit computation of these Tate cohomology groups when $G$ is ${\rm GL}_{n}$ and $\Pi$ is obtained as a base change lifting of a depth-zero cuspidal representation of ${\rm GL}_{n}(F)$. The primary novelty from our previous work is that we treat the case where $\Pi$ is possibly non-cuspidal. We also study the ${\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})$-Tate cohomology groups of the $-\ell$ reduction of the unipotent cuspidal representation of ${\rm Sp}_{4}(\mathbb{F}_{q^{\ell}})$.
1. Introduction
Let $G$ be a connected reductive algebraic group defined over a $p$-adic field $F$. Let $\ell$ be a prime number distinct from $p$. Let $\sigma$ be an order $\ell$ automorphism of $G$, defined over $F$, and let $\Gamma$ be the cyclic group generated by $\sigma$. We denote by $G^{\sigma}$ the connected component of the subgroup of $\sigma$-fixed points of $G$. Let $\Pi$ be an $\ell$-modular smooth irreducible representation of $G(F)$ such that $\Pi$ extends to a representation of $G(F)\rtimes\Gamma$. Treumann and Venkatesh, in their work on $-\ell$ functoriality ([TV16, Conjectures 6.3]), proposed some remarkable conjectures relating the $\ell$-modular irreducible smooth representations of the $p$-adic groups $G(F)$ and $G^{\sigma}(F)$. Their conjectures predict that the Tate cohomology groups $\widehat{H}^{i}(\Gamma,\Pi)$ are finite length representations of $G^{\sigma}(F)$, for $i\in\{0,1\}$, and they also proposed that the sub-quotients of $\widehat{H}^{i}(\Gamma,\Pi)$ must be related to $\Pi$ via $\ell$-modular functoriality in the sense of Langlands, and they called it the linkage principle. In our previous work [ND24], we proved that $\widehat{H}^{i}(\Gamma,\Pi)$ is a finite length representation for the case where $G={\rm Res}_{E/F}{\rm GL}_{n}$ and $\sigma$ is the automorphism of $G$ induced by a generator of the cyclic group ${\rm Gal}(E/F)$ of order $\ell$, and we showed that the linkage principle holds when $\ell$ does not divide the pro-order of ${\rm GL}_{n-1}(F)$.
In this article, we first prove the finiteness part of Treumann–Venkatesh conjecture for the case where $\sigma$ is induced by a Galois automorphism, i.e., the case where $G$ is ${\rm Res}_{E/F}H$ and $H$ is a connected reductive algebraic group defined over $F$, the extension $E/F$ is a cyclic Galois extension of degree $\ell$, and $\sigma$ is induced by a generator of the Galois group ${\rm Gal}(E/F)$. The main technique is to control the Jacquet modules and twisted Jacquet modules of the Tate cohomology groups. We also use the fact that subrepresentations of finitely generated $\ell$-modular representations of a $p$-adic group are finitely generated; this is a deep result proved recently in [DHKM24]. Let $\Pi$ be the base change lift to ${\rm GL}_{n}(E)$ of an integral cuspidal irreducible representation $\pi$ of ${\rm GL}_{n}(F)$ defined over the maximal unramified extension $\mathcal{K}$ of $\mathbb{Q}_{\ell}$. We further show that $\widehat{H}^{0}({\rm Gal}(E/F),\mathcal{L})$ is irreducible and a cuspidal representation, for any ${\rm GL}_{n}(E)\rtimes{\rm Gal}(E/F)$-stable $\mathbb{Z}_{\ell}^{\rm un}$-lattice $\mathcal{L}$ in $\Pi$. In the depth-zero case, using local methods, we obtain the precise description of Tate cohomology of a base change lift. These results on depth-zero representations are obtained by studying the Tate cohomology of invariant lattices in Shintani lifts for ${\rm GL}_{n}$ over finite fields.
In two papers [Fen23] and [Fen24], Feng showed that, under mild restrictive hypothesis on the groups involved, the Genestier–Lafforgue correspondence and the Fargues–Scholze correspondence are compatible with the linkage principle. However, the explicit computation of Tate cohomology groups, in particular, admissibility and the non-vanishing of Tate cohomology $\widehat{H}^{i}(\Gamma,\Pi)$, genericity, are not well understood. We answer some of these questions in this article. In their paper [HV25, Section 5.0.3], Henniart and Vignéras verified Treumann and Venkatesh conjectures for $G(F)={\rm GL}_{2}(F)$ and $G^{\sigma}(F)={\rm SL}_{2}(F)$ with $\ell=2$.
The following are our principal results. The first result of this article is to show that the Tate cohomology groups are finitely generated representations. We prove the following result.
Theorem 1.0.1.
Let $G$ be a connected reductive algebraic group over a non-Archimedean local field $F$. Let $\Gamma\subset{\rm Aut}(G)(F)$ be a subgroup of automorphisms of order $\ell$. Let $(\Pi,V)$ be a finite length $\ell$-modular representation of $G(F)\rtimes\Gamma$. Then the Tate cohomology $\widehat{H}^{i}(\Gamma,\Pi)$ is a finitely generated representation of $G^{\sigma}(F)$.
The above theorem is discussed in Section 2. The proof of Theorem 1.0.1 uses the $\ell$-modular version of Schneider and Stuhler resolutions ([SS97, Theorem II.3.1]) of smooth irreducible representations as in Vignéras’s work [Vig97, Proposition 2.6]. Note that we do not assume anything about $\Gamma$ for the above theorem. However, for the next part on the finite length of the Tate cohomology groups, we restrict to the Galois automorphism case. In what follows, we use the hypothesis and notations of the above theorem.
To prove the admissibility of Tate cohomology groups, we prove that the Jacquet modules of $\widehat{H}^{i}(\Gamma,\Pi)$ are controlled by the Tate cohomology of Jacquet modules of $\Pi$ with respect to $\sigma$-stable parabolic subgroups. This is discussed in Section 3. Let $P$ be a $\sigma$-stable parabolic subgroup of $G$. Then the connected component of the subgroup of $\sigma$-fixed points, denoted by $P^{\sigma}$, is a parabolic subgroup of $G^{\sigma}$. Let $U$ be the unipotent radical of $P$ and let $N$ be the unipotent radical of $P^{\sigma}$. Let $\Psi$ be a character of $U(F)$. We will show that the natural map of twisted Jacquet modules
$\widehat{H}^{i}(\Gamma,\Pi)_{N(F),\Psi}\rightarrow\widehat{H}^{i}(\Gamma,\Pi_{U(F),\Psi}),$ (1.1)
is an injection. This article, in essence, is about various applications of the above injective map. First of which is the following theorem.
Theorem 1.0.2.
Let $G$ be a connected reductive group defined over a non-Archimedean local field $F$, and let $E/F$ be a cyclic Galois extension of degree $\ell$. Let $\Pi$ be a finite length $\ell$-modular representation of $G(E)\rtimes{\rm Gal}(E/F)$. Then the Tate cohomology group $\widehat{H}^{i}({\rm Gal}(E/F),\Pi)$ is a finite length $\ell$-modular representation of $G(F)$.
We prove the above theorem using induction on the dimension of $G$. In the induction argument, we need to control the action of the centre of $G^{\sigma}$ along with the exponents in the Jacquet modules with respect to the parabolic subgroups of $G^{\sigma}$. For this reason we restrict to the case where $\sigma$ is induced by a Galois automorphism. In the Galois automorphism case, for a Levi subgroup $M$ defined over $F$, the action of the centre of $M(F)$ on the ${\rm Gal}(E/F)$-Tate cohomology of the corresponding Jacquet modules of $\Pi$ is via the restriction of the action of the centre of $M(E)$. In other words, we are using the fact that $Z(M(E))\cap G(F)=Z(M(F))$. In the general case, i.e., when $\sigma$ is not necessarily induced by a Galois automorphism, we note that $Z(G^{\sigma})$ can be quite large compared to $Z(G)^{\sigma}$. For instance, $\sigma$ could be an inner automorphism induced by a regular element of order $\ell$, and in this case $G^{\sigma}$ is a torus.
For the next result, proved in Section 4, we assume that $G$ is ${\rm Res}_{E/F}{\rm GL}_{n}$, and $\sigma$ is an automorphism of $G$ induced by a generator of the Galois group ${\rm Gal}(E/F)$. Let $\mathcal{K}$ be the maximal unramified extension of $\mathbb{Q}_{\ell}$ and $\Lambda$ be the ring of integers of $\mathcal{K}$. Let $\Pi$ be an absolutely irreducible, integral representation of ${\rm GL}_{n}(E)$, defined over $\mathcal{K}$ such that $\Pi^{\sigma}\simeq\Pi$. Let $\pi$ be an irreducible $\ell$-adic integral cuspidal representation of ${\rm GL}_{n}(F)$ such that $\Pi\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}$ is the base change lifting of $\pi$. Note that $\Pi$ is not necessarily a cuspidal representation. We show the following result on any $\Lambda[{\rm GL}_{n}(E)\rtimes{\rm Gal}(E/F)]$ stable lattice in $\Pi$.
Theorem 1.0.3.
With the Hypothesis on $\Pi$ and $\pi$ as above, let $\mathcal{L}$ be a $\Lambda[{\rm GL}_{n}(E)\rtimes{\rm Gal}(E/F)]$ stable lattice in $\Pi$. Then, the Tate cohomology $\widehat{H}^{0}({\rm Gal}(E/F),\mathcal{L})$ is an $\ell$-modular irreducible cuspidal representation of ${\rm GL}_{n}(F)$ and $\widehat{H}^{1}({\rm Gal}(E/F),\mathcal{L})$ is trivial.
In the proof of the above result, we first prove that the ${\rm Gal}(E/F)$-Tate cohomology of a $\Lambda[{\rm GL}_{n}(E)\rtimes{\rm Gal}(E/F)]$ stable lattice in $\Pi$ has a unique generic sub-quotient. In fact, this genericity result holds in greater generality. Let $G$ be a split connected reductive group defined over $F$. Let $\Pi$ be a smooth irreducible $\ell$-modular representation of $G(E)$ such that $\Pi$ is stable under the action of ${\rm Gal}(E/F)$. Let $B$ be a Borel subgroup of $G$ defined over $F$, and let $U$ be the unipotent radical of $B$. Let $\psi:U(E)\rightarrow\overline{\mathbb{F}}_{\ell}^{\times}$ be a non-degenerate, ${\rm Gal}(E/F)$ invariant character such that $\psi$ restricts to a non-degenerate character on $U(F)$. If $\Pi$ is $(U(E),\psi)$-generic, then we show that $\widehat{H}^{0}({\rm Gal}(E/F),\Pi)$ has a unique $(U(F),\psi)$-generic sub-quotient (Theorem 4.0.1). Then to prove Theorem 1.0.3, we use the injection (1.1) and prove the right-hand side of the equation (1.1) is trivial by using the fact that the non-trivial Jacquet modules of a non-cuspidal base change lifting $\Pi$ are multiplicity free and their Jordan Holder factors are not stable under ${\rm Gal}(E/F)$.
In Section 5, we then study the Tate cohomology of invariant lattices in representations obtained as Shintani liftings (to ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$) of representations of finite group ${\rm GL}_{n}(\mathbb{F}_{q})$. The analogue of Theorem 1.0.3 is valid for $\Pi$ which is obtained as a Shintani lift to ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$ of a cuspidal representation of ${\rm GL}_{n}(\mathbb{F}_{q})$; we use the Deligne–Lusztig character formula to show that, for any $\Lambda[{\rm GL}_{n}(\mathbb{F}_{q^{\ell}})\rtimes{\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})]$ stable lattice $\mathcal{L}$ in $\Pi$, the space $\widehat{H}^{0}({\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}),\mathcal{L})$ is isomorphic to the Frobenius twist of the $-\ell$ reduction of $\pi$ (Theorem 5.4.3).
We end this article by considering the example of $\theta_{10}$ in Section 7, the unipotent cuspidal irreducible representation of the symplectic group ${\rm Sp}_{4}(\mathbb{F}_{q})$. Let $\pi_{n}$ be the unique cuspidal unipotent representation of ${\rm Sp}_{4}(\mathbb{F}_{q^{n}})$. Note that the $-\ell$ reduction of $\pi_{n}$ is irreducible and it is denoted by $r_{\ell}(\pi_{n})$. For a vector space $W$ over the field $\overline{\mathbb{F}}_{\ell}$, we denote by $W^{(\ell)}$ the $\overline{\mathbb{F}}_{\ell}$-vector space where the scalar action is twisted by the Frobenius automorphism of the field $\overline{\mathbb{F}}_{\ell}$. We have
Theorem 1.0.4.
Let $\ell$ be an odd prime number. Let $\pi_{n}$ be the unipotent cuspidal representation of ${\rm Sp}_{4}(\mathbb{F}_{q^{n}})$. Then, we have the following isomorphism of ${\rm Sp}_{4}(\mathbb{F}_{q})$ representations:
$\widehat{H}^{0}({\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}),r_{\ell}(\pi_{\ell}))\simeq r_{\ell}(\pi_{1})^{(\ell)}.$
Moreover, the Tate cohomology $\widehat{H}^{1}({\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}),r_{\ell}(\pi_{\ell}))$ is trivial.
The above theorem seems to suggest that for all classical groups $G$ admitting a cuspidal unipotent representation, the zeroth Tate cohomology group for the action of ${\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})$ on the $-\ell$ reduction of a cuspidal unipotent representation of $G(\mathbb{F}_{q^{\ell}})$ is isomorphic to the Frobenius twist of the $-\ell$ reduction of the cuspidal unipotent representation of $G(\mathbb{F}_{q})$.
Notations
  1. (1)
    We use the standard notations attached to any non-Archimedean local field $F$: the ring of integers of $F$ is $\mathfrak{o}_{F}$, the maximal ideal of $\mathfrak{o}_{F}$ is $\mathfrak{p}_{F}$, and the residue field $\mathfrak{o}_{F}/\mathfrak{p}_{F}$ is $k_{F}$. The field $k_{F}$ has $q_{F}$ elements and $q_{F}$ is a power of a prime $p$.
  2. (2)
    Let $G$ be a connected reductive algebraic group defined over $F$. Let $\sigma\in{\rm Aut}(G)$, defined over $F$, be an element of prime order $\ell$ and $\ell\neq p$. Let $\Gamma$ be the group generated by $\sigma$. Let $G^{\sigma}$ be the connected component of the fixed points of $\sigma$, and it is assumed to be reductive.
  3. (3)
    A representation $(\pi,V)$ of $G(F)$ is said to be smooth if every vector $v\in V$ is fixed by a compact open subgroup of $G(F)$. Let $P$ be an $F$-rational parabolic subgroup of $G$ and the unipotent radical of $P$ is denoted by $U$. Let $M$ be a Levi subgroup of $P$ defined over $F$. For a smooth representation $(\tau,W)$ of $M(F)$, we denote by $i_{P(F)}^{G(F)}(\tau)$ the normalised parabolic induction of $\tau$ to $G(F)$. We use the notation $r_{U(F)}(\pi)$ for the normalised Jacquet module of a smooth representation $(\pi,V)$ of $G(F)$.
  4. (4)
    In this article, a smooth representation $(\pi,V)$ of $G(F)$ is called cuspidal if $r_{U(F)}(\pi)=0$, where $U$ is the unipotent radical of any proper $F$-rational parabolic subgroup $P$ of $G$ For a closed subgroup $H_{1}$ of a reductive $p$-adic group $H$, we will denote by ${\rm Ind}_{H_{1}}^{H}$ and ${\rm ind}_{H_{1}}^{H}$ the smooth induction functor and the compact induction functor, respectively.
2. Tate cohomology of smooth representations
Let $\Pi$ be a smooth finite length representation of $\overline{\mathbb{F}}_{\ell}[G(F)\rtimes\Gamma]$. We prove that the Tate cohomology groups $\widehat{H}^{i}(\Gamma,\Pi)$ are finitely generated representations of $G^{\sigma}(F)$. We use Schneider–Stuhler resolutions and their $-\ell$ version by Vignéras for proving this finiteness statement.
2.1.
Fix an algebraic closure $\overline{\mathbb{Q}}_{\ell}$ of $\mathbb{Q}_{\ell}$ and let $\mathcal{K}$ be the maximal unramified extension of $\mathbb{Q}_{\ell}$ in $\overline{\mathbb{Q}}_{\ell}$. Let $\Lambda$ be the ring of integers of $\mathcal{K}$. All $\Lambda[G(F)]$ modules are assumed to be smooth, i.e., every element is fixed by an open subgroup of $G(F)$. Let $M$ be a $\Lambda[G(F)\rtimes\Gamma]$ module. Let ${\rm Nr}$ be the norm element $\sum_{g\in\Gamma}g\in\Lambda[\Gamma]$. Consider the chain complex of $\Lambda[\Gamma]$ modules
$\cdots\longrightarrow M\xrightarrow{{\rm id}-\sigma}M\xrightarrow{{\rm Nr}}M\xrightarrow{{\rm id}-\sigma}M\xrightarrow{{\rm Nr}}M\longrightarrow\cdots$
The Tate cohomology groups of $M$, with respect to the action of $\Gamma$, are 222-periodic and we have
$\widehat{H}^{0}(\Gamma,M):=\dfrac{{\rm Ker}({\rm id}-\sigma)}{{\rm Img}({\rm Nr})},\,\,\,\widehat{H}^{1}(\Gamma,M):=\dfrac{{\rm Ker}({\rm Nr})}{{\rm Img}({\rm id}-\sigma)}.$
The action of $G(F)$ on $M$ induces an action of $G^{\sigma}(F)$ on the Tate cohomology groups $\widehat{H}^{i}(\Gamma,M)$, and we get two $\ell$-modular smooth representations $\widehat{H}^{i}(\Gamma,M)$, for $i\in\{0,1\}$, of the fixed point subgroup $G^{\sigma}(F)$. Similar cohomology groups can be defined when $F$ is a finite field, and we will assume the easy modifications in the definitions.
2.2.
Let $X$ be a Hausdorff, locally compact, and totally disconnected topological space (an $\ell$-space in the sense of [BZ76, Section 1.1]). Let $G$ be a locally profinite group with an order $\ell$ automorphism $\sigma$, generating a group $\Gamma$ of automorphisms of $G$. Assume that $G\rtimes\Gamma$ acts continuously on $X$. Let $\mathcal{F}$ be a $G\rtimes\Gamma$ equivariant sheaf on $X$. The natural restriction map
$\widehat{H}^{i}(\Gamma,H^{0}(X,\mathcal{F}))\rightarrow H^{0}(X^{\sigma},\widehat{H}^{i}(\Gamma,\mathcal{F}_{|X^{\sigma}}))$
is an isomorphism, for $i\in\{0,1\}$ (see [TV16, Proposition 3.3]). Hear, $H^{0}(X,\mathcal{F})$ is the space of global sections of a sheaf $\mathcal{F}$ on $X$. Let $H$ be a closed subgroup of $G$ such that $\sigma(H)=H$. Let $V$ be a $\Lambda[H\rtimes\Gamma]$ module. This data defines a $G\rtimes\Gamma$ equivariant sheaf $\mathcal{F}_{V}$ on $G/H$ whose sections on an open set $U$ of $G/H$ are the sections of the vector bundle $G\times_{H}V\rightarrow G/H$ restricted to $U$. The global sections of the sheaf $\mathcal{F}_{V}$ is the induced representation ${\rm ind}_{H}^{G}V$, and the action of $\Gamma$ is given by
$(\sigma.f)(g)=\sigma\big{(}f(\sigma^{-1}(g))\big{)},$
for all $g\in G$ and $f\in{\rm ind}_{H}^{G}V$. Assume that $(G/H)^{\sigma}=G^{\sigma}/H^{\sigma}$. Then, the restriction to $G^{\sigma}$ map induces a $G^{\sigma}$ equivariant isomorphism
$\widehat{H}^{i}(\Gamma,{\rm ind}_{H}^{G}V)\rightarrow{\rm ind}_{H^{\sigma}}^{G^{\sigma}}(\widehat{H}^{i}(\Gamma,V)),\,\,\,f\mapsto{\rm res}_{G^{\sigma}}(f)$ (2.1)
2.3.
Let $(\Pi,V)$ be a smooth finite length representation of $\overline{\mathbb{F}}_{\ell}[G(F)\rtimes\Gamma]$. We show that $\widehat{H}^{i}(\Gamma,\Pi)$ is a finitely generated $G^{\sigma}(F)$ representation. We use the $-\ell$ version of Schneider and Stuhler resolutions ([SS97, Theorem II.3.1]) of smooth irreducible representations as discussed in Vignéras ([Vig97, Proposition 2.6]). Let $\mathfrak{X}$ be the extended Bruhat–Tits building of a connected reductive algebraic group $G$ defined over a non-Archimedean local field $F$. Let $d$ be the $F$-rank of $G$. For $0\leq k\leq d$, let $\mathfrak{X}_{k}$ be the set of $k$-dimensional facets of $\mathfrak{X}$, and let $\mathfrak{X}_{(k)}$ be the set of oriented $k$-facets which are tuples of the form $(\mathcal{F},c_{\mathcal{F}})$, where $\mathcal{F}\in\mathfrak{X}_{k}$ and $c_{\mathcal{F}}$ is an orientation of $\mathcal{F}$. Let $e\geq 0$. For a facet $\mathcal{F}$ of $\mathfrak{X}$, Schneider and Stuhler associate a compact open subgroup $U_{\mathcal{F}}^{(e)}$, and let $\underline{V}$ be the coefficient system defined by associating every facet $\mathcal{F}$ to $V^{U^{e}_{\mathcal{F}}}$. Let $C(\mathfrak{X}_{(k)},\underline{V})$ be the set of oriented cellular $k$-chains on $\mathfrak{X}_{(k)}$. Then we have a $G$ equivariant resolution of the representation
$0\rightarrow C(\mathfrak{X}_{(d)},\underline{V})\xrightarrow{\partial_{d-1}}C(\mathfrak{X}_{(d-1)},\underline{V})\xrightarrow{\partial_{d-2}}\cdots\rightarrow C(\mathfrak{X}_{(1)},\underline{V})\xrightarrow{\partial_{0}}C(\mathfrak{X}_{(0)},\underline{V})\xrightarrow{\epsilon}V\rightarrow 0$ (2.2)
for some $e\geq 0$. We refer to [SS97, Section II.2] for details. Note the automorphism $\sigma$ acts on $\mathfrak{X}$ simplicially, i.e., it preserves the sets $\mathfrak{X}_{d}$ and $\mathfrak{X}_{(d)}$.
Lemma 2.3.1.
For $0\leq k\leq d$, the Tate cohomology groups $\widehat{H}^{i}(\Gamma,C(\mathfrak{X}_{(k)},\underline{V}))$ are finitely generated $\overline{\mathbb{F}}_{\ell}[G^{\sigma}(F)]$ modules.
Proof.
Consider the set $\mathfrak{X}_{(k)}$ with discrete topology, and the action of the group $G(F)$ on $\mathfrak{X}_{(k)}$ is continuous. Also note that the sheaf $\underline{V}$ is $G$ and $\Gamma$ equivariant. Since $V$ is an admissible representation of $\overline{\mathbb{F}}_{\ell}[G(F)]$, the stalks of the sheaf $\underline{V}$ are all finite dimensional vector spaces over $\overline{\mathbb{F}}_{\ell}$. Using [TV16, Proposition 3.3], we get that the restriction map
$\widehat{H}^{i}(\Gamma,C(\mathfrak{X}_{(k)},\underline{V}))\rightarrow C(\mathfrak{X}_{(k)}^{\sigma},\widehat{H}^{i}(\Gamma,\underline{V}))$
is an isomorphism. Let $\mathcal{Y}$ be the Bruhat–Tits building of $G^{\sigma}$, then $\mathfrak{X}^{\sigma}_{k}$ is a subset of facets of $\mathcal{Y}$. Thus, there are only finitely many orbits for the action of $G^{\sigma}$ on $\mathfrak{X}^{\sigma}_{(k)}$. Since the module $\widehat{H}^{i}(\Gamma,V^{U_{F}^{e}})$ is finite, we get that the space $C(\mathfrak{X}_{(k)}^{\sigma},\widehat{H}^{i}(\Gamma,\underline{V}))$ is a finitely generated representation of $\overline{\mathbb{F}}_{\ell}[G^{\sigma}(F)]$. This shows the lemma. ∎
In what follows, we need the fact that the submodules of finitely generated smooth $\overline{\mathbb{F}}_{\ell}[H(F)]$ representations are again finitely generated. This is a fundamental result in the theory of smooth complex representations due to Bernstein and in the $-\ell$ case it follows from the deep results of the paper [DHKM24, Corollary 1.3] seen with the work of Dat ([Dat09, Corollary 4.5]).
Lemma 2.3.2.
Let $(\Pi,V)$ be an irreducible smooth representation of $\overline{\mathbb{F}}_{\ell}[G(F)]$. Assume that $(\Pi,V)$ extends as a representation of $\overline{\mathbb{F}}_{\ell}[G(F)\rtimes\Gamma]$. Then the Tate cohomology groups $\widehat{H}^{i}(\Gamma,V)$ are finitely generated representations of $\overline{\mathbb{F}}_{\ell}[G^{\sigma}(F)]$.
Proof.
Note that the resolution (2.2) is $G$ and $\Gamma$ equivariant, and the Tate cohomology groups
[…]
are finitely generated representations of $\overline{\mathbb{F}}_{\ell}[G^{\sigma}(F)]$ by Lemma 2.3.1. For each $2\leq k\leq d$, consider the short exact sequence of $\overline{\mathbb{F}}_{\ell}[G(F)\rtimes\Gamma]$ modules
$0\rightarrow{\rm Im}(\partial_{k-1})\rightarrow C(\mathfrak{X}_{(k-1)},\underline{V})\rightarrow{\rm Ker}(\partial_{k-2})\rightarrow 0$
Then the corresponding long exact sequence of Tate cohomology groups gives the following hexagonal exact sequence
${\widehat{H}^{0}(\Gamma,{\rm Im}(\partial_{k-1}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\widehat{H}^{0}(\Gamma,C(\mathfrak{X}_{(k-1)},\underline{V}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\widehat{H}^{0}(\Gamma,{\rm Ker}(\partial_{k-2}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\widehat{H}^{1}(\Gamma,{\rm Ker}(\partial_{k-2}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\widehat{H}^{1}(\Gamma,C(\mathfrak{X}_{(k-1)},\underline{V}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\widehat{H}^{1}(\Gamma,{\rm Im}(\partial_{k-1}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
which we denote by $S(k)$. Note that ${\rm Im}(\partial_{d-1})$ is identified with $C(\mathfrak{X}_{(d)},\underline{V})$. Since both the Tate cohomology $\widehat{H}^{i}(\Gamma,C(\mathfrak{X}_{(d)},\underline{V}))$ and $\widehat{H}^{i}(\Gamma,C(\mathfrak{X}_{(d-1)},\underline{V}))$ are finitely generated, it follows from the diagram $S(d)$ that $\widehat{H}^{i}(\Gamma,{\rm Ker}(\partial_{d-2}))$ is also finitely generated. A similar hexagonal diagram of Tate cohomology corresponding to the following short exact sequence of $G(F)\rtimes\Gamma$ modules
$0\rightarrow{\rm Ker}(\partial_{d-2})\rightarrow C(\mathfrak{X}_{(d-1)},\underline{V})\rightarrow{\rm Im}(\partial_{d-2})\rightarrow 0$
shows that $\widehat{H}^{i}(\Gamma,{\rm Im}(\partial_{d-2}))$ is finitely generated. Since $\widehat{H}^{i}(\Gamma,C(\mathfrak{X}_{(d-2)},\underline{V}))$ is finitely generated, it follows from the hexagonal diagram $S(d-1)$ that $\widehat{H}^{i}(\Gamma,{\rm Ker}(\partial_{d-3}))$ is finitely generated. Thus, following a finite number of inductive steps, the hexagonal diagram $S(2)$ shows that the Tate cohomology $\widehat{H}^{i}(\Gamma,{\rm Ker}(\partial_{0}))$ and $\widehat{H}^{i}(\Gamma,{\rm Im}(\partial_{0}))$ are finitely generated. Finally, it follows from the hexagonal diagram of Tate cohomology groups corresponding to the following short exact sequence of $\overline{\mathbb{F}}_{\ell}[G(F)\rtimes\Gamma]$ modules
$0\rightarrow{\rm Im}(\partial_{0})\rightarrow C(\mathfrak{X}_{(0)},\underline{V})\xrightarrow{\epsilon}V\rightarrow 0,$
that the Tate cohomology $\widehat{H}^{i}(\Gamma,V)$ is finitely generated as a representation of $\overline{\mathbb{F}}_{\ell}[G^{\sigma}(F)]$. ∎
Theorem 2.3.3.
Let $(\Pi,V)$ be a finite length representation of $\overline{\mathbb{F}}_{\ell}[G(F)\rtimes\Gamma]$. Then each Tate cohomology group $\widehat{H}^{i}(\Gamma,V)$ is a finitely generated representation of $\overline{\mathbb{F}}_{\ell}[G^{\sigma}(F)]$.
Proof.
We use induction on the length of $(\Pi,V)$ as an $\overline{\mathbb{F}}_{\ell}[G(F)]$ representation. If $(\Pi,V)$ is an irreducible $G(F)$ representation, then the theorem follows from Lemma 2.3.2. Let $W$ be an irreducible $G(F)$ subrepresentation of $V$. First consider the case when $W$ is $\Gamma$ stable. The long exact sequence of Tate cohomology associated with the following short exact sequence
$0\longrightarrow W\longrightarrow V\longrightarrow V/W\longrightarrow 0$
gives two exact sequences
$\widehat{H}^{0}(\Gamma,W)\longrightarrow\widehat{H}^{0}(\Gamma,V)\longrightarrow\widehat{H}^{0}(\Gamma,V/W)$
and
$\widehat{H}^{1}(\Gamma,W)\longrightarrow\widehat{H}^{1}(\Gamma,V)\longrightarrow\widehat{H}^{1}(\Gamma,V/W).$
Since $W$ is irreducible, each Tate cohomology $\widehat{H}^{i}(\Gamma,W)$ is a finitely generated representation of $\overline{\mathbb{F}}_{\ell}[G^{\sigma}(F)]$ by Lemma 2.3.2. Also note that the length of the $G(F)$ representation $V/W$ is strictly less than the length of $V$, inductively we may assume that the Tate cohomology groups $\widehat{H}^{i}(\Gamma,V/W)$ are finitely generated representations of $G^{\sigma}(F)$. Now, it follows from the above exact sequences that each $\widehat{H}^{i}(\Gamma,V)$ is a finitely generated representation of $\overline{\mathbb{F}}_{\ell}[G^{\sigma}(F)]$. Let us now consider the case where $W$ is not $\Gamma$ stable. Then
$W^{\prime}=\bigoplus_{i=0}^{\ell-1}W^{\sigma^{i}}$
is a $G(F)\rtimes\Gamma$ stable subspace of $V$, where the action of $G(F)$ on the space $W^{\sigma^{i}}$ is twisted by $\sigma^{i}$ for $0\leq i<\ell$. Note that $\widehat{H}^{i}(\Gamma,W^{\prime})$ is trivial, for $i\in\{0,1\}$. Now, the long exact sequence of Tate cohomology groups applied to the following short exact sequence of $\overline{\mathbb{F}}_{\ell}[G(F)\rtimes\Gamma]$ modules
$0\longrightarrow W^{\prime}\longrightarrow V\longrightarrow V/W^{\prime}\longrightarrow 0$
gives the following isomorphism of $G^{\sigma}(F)$ representations
$\widehat{H}^{i}(\Gamma,V)\rightarrow\widehat{H}^{i}(\Gamma,V/W^{\prime}),\ i\in\{0,1\}.$
Since the length of $V/W^{\prime}$ is strictly less than that of $V$, the $G^{\sigma}(F)$ representation $\widehat{H}^{i}(\Gamma,V/W^{\prime})$ is finitely generated by induction. Thus, we get that the Tate cohomology group $\widehat{H}^{i}(\Gamma,V)$ is a finitely generated representation of $G^{\sigma}(F)$. ∎
3. Finiteness of Tate cohomology, the Galois automorphisms case
In this section, we prove some results on the compatibility of Jacquet modules and twisted Jacquet modules with Tate cohomology groups. Some of these results were proved in [ND24, Section 6] for the group $G={\rm Res}_{E/F}{\rm GL}_{n}$, where $E/F$ is a Galois extension of prime degree $\ell$, and $\sigma$ is an automorphism induced by a generator of the Galois group ${\rm Gal}(E/F)$. The results in this section, although proved for groups defined over non-Archimedean local fields, are valid for finite groups of Lie type. For convenience, we will drop $\Gamma$ in the Tate cohomology groups $\widehat{H}^{i}(\Gamma,M)$ and simply denote these cohomology spaces by $\widehat{H}^{i}(M)$.
3.1. Twisted Jacquet modules
Let $(\pi,V)$ be a smooth $\Lambda[G(F)]$ module, where $\Lambda$ is the ring of integers of the maximal unramified extension of $\mathbb{Q}_{\ell}$ in $\overline{\mathbb{Q}}_{\ell}$. Let $P=MN$ be an $F$-rational parabolic subgroup of $G$, where $M$ is an $F$-rational Levi subgroup of $P$ and $N$ is the unipotent radical of $P$. Let $\Theta:N(F)\rightarrow\Lambda^{\ast}$ be a character. Let $V(N(F),\Theta)$ be the $\Lambda$-module generated by the following set of vectors
$\big{\{}\pi(n)v-\Theta(n)v:v\in V,\,n\in N(F)\big{\}}.$
The quotient space $V/V(N(F),\Theta)$ is called a twisted Jacquet module of $V$, which is denoted by $V_{N(F),\Theta}$. This space is characterised by the set of elements $v\in V$ for which there exists some compact open subgroup $\mathcal{N}_{v}$ of $N(F)$ such that
[…]
where $dn$ is a Haar measure on $N(F)$. The space $V_{N(F),\Theta}$ is a smooth representation of the $M(F)$ stabiliser of the character $\Theta$, to be denoted by $M(F)_{\Theta}$. In particular, when $\Theta=1$, the trivial character of $N(F)$, then the smooth $\Lambda[M(F)]$ module $V_{N(F),1}$ is called the Jacquet module of $V$. We also use the notation $r_{N(F)}(V)$ for $V_{N(F),1}$.
3.2. Integral representation and $-\ell$ reduction
A smooth $\ell$-adic representation $(\pi,V)$ of $G(F)$, i.e., $V$ is a vector space over $\overline{\mathbb{Q}}_{\ell}$, is called integral if $V$ is a finite length $G(F)$ representation and if there exists a $G(F)$ stable $\overline{\mathbb{Z}}_{\ell}$-lattice $\mathcal{L}$ in $V$. Suppose $(\pi,V)$ is defined over $\mathcal{K}$, i.e., there exists a $G(F)$ stable $\mathcal{K}$-vector subspace $V_{0}$ in $V$ such that $V_{0}\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}=V$, then there exists a $\Lambda[G(F)]$ stable lattice $\mathcal{L}_{0}$ in $V_{0}$ and $\mathcal{L}_{0}\otimes_{\Lambda}\overline{\mathbb{Z}}_{\ell}$ is a $G(F)$ stable $\overline{\mathbb{Z}}_{\ell}$-lattice in $V$. If $P$ is an $F$-rational parabolic subgroup of $G$ with an $F$-rational Levi subgroup $M$ and unipotent radical $N$, then it follows from [Dat05, Proposition 1.4] that $\mathcal{L}_{N(F)}$ is a $M(F)$ stable $\overline{\mathbb{Z}}_{\ell}$-lattice in $V_{N(F)}$. The natural action of $G(F)$ on the $\overline{\mathbb{F}}_{\ell}$-vector space ¯$\mathcal{L}\otimes_{\overline{\mathbb{Z}}_{\ell}}\overline{\mathbb{F}}_{\ell}$ gives an $\ell$-modular representation of $G(F)$, which depends on the choice of the lattice $\mathcal{L}$. Using [Vig96, Chapter II, 5.11.a and 5.11.b], the representation $(\pi,\mathcal{L}\otimes_{\overline{\mathbb{Z}}_{\ell}}\overline{\mathbb{F}}_{\ell})$ has finite length and its semi-simplification is independent of the choice of the $G(F)$ stable $\overline{\mathbb{Z}}_{\ell}$-lattice $\mathcal{L}$ in $V$. The semi-simplification of $(\pi,\mathcal{L}\otimes_{\overline{\mathbb{Z}}_{\ell}}\overline{\mathbb{F}}_{\ell})$, denoted by $r_{\ell}(\pi)$, is called the $-\ell$ reduction of $\pi$.
3.3.
Let $\sigma\in{\rm Aut}(G)$, defined over $F$, be an element of order $\ell$, where $\ell$ is a prime different from $p$. Let $\Gamma$ be the group generated by $\sigma$. Let $P$ be a $\sigma$-stable $F$-parabolic subgroup of $G$, and let $N$ be the unipotent radical of $P$. Note that $N$ is $\sigma$-stable and let $M$ be a $\sigma$-stable Levi subgroup $P$ (for the existence of $M$, see [KP23, Lemma 12.5.4] for a reference). Given any $F$-parabolic subgroup $Q$ of $G^{\sigma}$, there exists a $\sigma$-stable $F$-parabolic subgroup $P$ of $G$ such that $Q=P^{\sigma}$ (see [KP23, Lemma 12.5.4] for a reference). Note that $U=N^{\sigma}$ is the unipotent radical of $Q$ and $M^{\sigma}$ is a Levi subgroup of $Q$. Let $\Theta:N(F)\rightarrow\Lambda^{\ast}$ be a character such that $\Theta^{\sigma}=\Theta$.
Lemma 3.3.1.
Let $(\pi,\mathcal{L})$ be a smooth $\Lambda[G(F)\rtimes\Gamma]$ module. For each $i\in\{0,1\}$, the image of the natural map $\widehat{H}^{i}(\mathcal{L}(N(F),\Theta))\rightarrow\widehat{H}^{i}(\mathcal{L})$ is equal to $\widehat{H}^{i}(\mathcal{L})(N^{\sigma}(F),\overline{\Theta})$, where $\overline{\Theta}$ is the the $-\ell$ reduction of $\Theta$.
Proof.
For $i\in\{0,1\}$, let $\varphi_{i}$ be the natural map
$\widehat{H}^{i}(\mathcal{L}(N(F),\Theta))\longrightarrow\widehat{H}^{i}(\mathcal{L}).$
We first consider the case $i=0$. Let $v$ be in the image of $\varphi_{0}$, and let $\tilde{v}$ be a lift of $v$ in $\mathcal{L}(N(F),\Theta)^{\sigma}$. There exists a compact open subgroup $\mathcal{N}\subseteq N(F)$ such that
[…]
Since $(\pi,\mathcal{L})$ is smooth, there exists a compact open subgroup $\mathcal{N}^{\prime}\subseteq\mathcal{N}$ of finite index such that $\mathcal{N^{\prime}}$ fixes the vector $\tilde{v}$ and $\Theta$ is trivial on $\mathcal{N}^{\prime}$. Then
[…]
Since $N(F)$ admits a filtration of $\Gamma$ stable compact open subgroups, we assume that both $\mathcal{N}$ and $\mathcal{N}^{\prime}$ are $\Gamma$-stable. If $X$ denotes the coset space $\mathcal{N}/\mathcal{N}^{\prime}$, then we have
$\sum_{x\in X}\Theta^{-1}(x)\,\pi(x)\tilde{v}=\sum_{y\in X^{\sigma}}\Theta^{-1}(y)\,\pi(y)\tilde{v}\ +\sum_{z\in{X\setminus X^{\sigma}}}\Theta^{-1}(z)\,\pi(z)\tilde{v}.$ (3.2)
Since the action of $\Gamma$ on $X\setminus X^{\sigma}$ is free, there exists a subset $\mathcal{U}\subseteq X\setminus X^{\sigma}$ such that $X\setminus X^{\sigma}$ is the disjoint union of $\sigma^{i}\mathcal{U}$ for $1\leq i\leq\ell$. As $\tilde{v}$ is a $\sigma$-fixed vector, we get
$\sum_{z\in{X\setminus X^{\sigma}}}\Theta^{-1}(z)\,\pi(z)\tilde{v}=\sum_{i=1}^{\ell}\sum_{u\in\mathcal{U}}\Theta^{-1}(\sigma^{i}u)\,\pi(\sigma^{i}u)\tilde{v}={\rm Nr}\big{(}\sum_{u\in\mathcal{U}}\Theta^{-1}(u)\,\pi(u)\tilde{v}\big{)},$
where ${\rm Nr}=1+\sigma+\cdots+\sigma^{\ell-1}\in\Lambda[\Gamma]$. This shows that
$\sum_{z\in{X\setminus X^{\sigma}}}\overline{\Theta}^{-1}(z)\,\pi(z)v=0$
in $\widehat{H}^{0}(\mathcal{L})$. Then it follows from (3.1) and (3.2) that
$\sum_{y\in X^{\sigma}}\overline{\Theta}^{-1}(y)\,\pi(y)v=0.$
This implies that the vector $v$ belongs to the Tate cohomology space $\widehat{H}^{0}(\mathcal{L})(N^{\sigma}(F),\overline{\Theta})$. Conversely, let $w$ be an element of $\widehat{H}^{0}(\mathcal{L})(N^{\sigma}(F),\overline{\Theta})$. Let $\mathcal{N}_{\sigma}$ be the compact open subgroup of $N^{\sigma}(F)$ with
[…]
Since $\pi$ is smooth, there exists a compact open subgroup $\mathcal{N}_{\sigma}^{\prime}\subseteq\mathcal{N}_{\sigma}$ of finite index such that
[…]
Let $\tilde{w}$ be a lift of $w$ in $\mathcal{L}^{\sigma}$. Consider the element
$\tilde{w}_{1}=\tilde{w}-\frac{1}{\lvert\mathcal{N}_{\sigma}/\mathcal{N}_{\sigma}^{\prime}\rvert}\sum_{n\in\mathcal{N}_{\sigma}/\mathcal{N}_{\sigma}^{\prime}}\Theta^{-1}(n)\,\pi(n)\tilde{w}.$
Then $\tilde{w}_{1}$ belongs to $\mathcal{L}(N(F),\Theta)^{\sigma}$ and $\varphi_{0}(\tilde{w}_{1})=w$. Thus, we get the lemma for $i=0$.
Now, consider the case $i=1$. Let $v$ be an element in the image of $\varphi_{1}$. Let $\tilde{v}$ be a lift of $v$ in ${\rm Ker}({\rm Nr}_{|\mathcal{L}(N(F),\Theta)})$. Then there exists $\Gamma$ stable compact open subgroups $\mathcal{N}^{\prime}\subset\mathcal{N}$ such that $\mathcal{N}^{\prime}$ fixes the vector $\tilde{v}$, the character $\Theta$ is trivial on $\mathcal{N}^{\prime}$, and
$\sum_{x\in X}\Theta^{-1}(x)\,\pi(x)\tilde{v}=0,$
where $X$ denotes the quotient space $\mathcal{N}/\mathcal{N}^{\prime}$. Since the $\Gamma$ action on $X\setminus X^{\sigma}$ is free, there exists a subset $\mathcal{V}\subseteq X\setminus X^{\sigma}$ such that $X\setminus X^{\sigma}=\coprod_{i=0}^{\ell-1}\sigma^{i}\mathcal{V}$. Then we have the identity
$\sum_{x\in X}\Theta^{-1}(X)\,\pi(x)\tilde{v}=\sum_{x\in\mathcal{V}}\Theta^{-1}(x)\,\pi(x){\rm Nr}(\tilde{v})+\sum_{x\in X^{\sigma}}\Theta^{-1}(x)\,\pi(x)\tilde{v}.$
Thus, we get that $\sum_{x\in X^{\sigma}}\overline{\Theta}^{-1}(x)\,\pi(x)v$ is trivial in $\widehat{H}^{1}(\mathcal{L})$, and this implies that $v\in\widehat{H}^{1}(\mathcal{L})(N^{\sigma}(F),\overline{\Theta})$. Conversely, let $v$ be an element in $\widehat{H}^{1}(\mathcal{L})(N^{\sigma}(F),\overline{\Theta})$. Let $\tilde{v}\in{\rm Ker}({\rm Nr})$ be a lift of $v$. Consider the element
$\tilde{v}_{1}=\tilde{v}-\dfrac{1}{|\mathcal{N}_{\sigma}/\mathcal{N}_{\sigma}^{\prime}|}\sum_{n\in\mathcal{N}_{\sigma}/\mathcal{N}_{\sigma}^{\prime}}\Theta^{-1}(n)\,\pi(n)\tilde{v}.$
Note that $\tilde{v}_{1}\in\mathcal{L}(N(F),\Theta)$ and ${\rm Nr}(\tilde{v}_{1})=0$ as ${\rm Nr}$ commutes with the action of $N^{\sigma}(F)$. This completes the proof. ∎
Using the above lemma, we get the following crucial result on Jacquet modules of the Tate cohomology groups.
Theorem 3.3.2.
Let $P$ be a $\sigma$ stable $F$-parabolic subgroup of $G$, and let $N$ be the unipotent radical of $P$. Let $\mathcal{L}$ be a smooth $\Lambda[G(F)\rtimes\Gamma]$ module. Then the natural map $\widehat{H}^{i}(\mathcal{L})_{N^{\sigma}(F),\overline{\Theta}}\rightarrow\widehat{H}^{i}(\mathcal{L}_{N(F),\Theta})$ is injective.
Proof.
Consider the short exact sequence of $\Lambda[\Gamma]$ modules
$0\longrightarrow\mathcal{L}(N(F),\Theta)\longrightarrow\mathcal{L}\longrightarrow\mathcal{L}_{N(F),\Theta}\longrightarrow 0$
The long exact sequence of Tate cohomology corresponding to the above short exact sequence gives the following eact sequence:
$\widehat{H}^{0}(\mathcal{L}(N(F),\Theta))\xrightarrow{\varphi_{0}}\widehat{H}^{0}(\mathcal{L})\rightarrow\widehat{H}^{0}(\mathcal{L}_{N(F),\Theta})\rightarrow\widehat{H}^{1}(\mathcal{L}(N(F),\Theta))\xrightarrow{\varphi_{1}}\widehat{H}^{1}(\mathcal{L})\rightarrow\widehat{H}^{1}(\mathcal{L}_{N(F),\Theta}).$
Using Lemma 3.3.1, the image of $\varphi_{i}$ is equal to $\widehat{H}^{i}(\mathcal{L})(N^{\sigma}(F),\overline{\Theta})$. Hence, we get the theorem. ∎
As a corollary, we get the following result.
Corollary 3.3.3.
Let $G$ be a connected reductive algebraic group defined over $F$, and let $\sigma$ be an order $\ell$ automorphism of $G$, where $\ell$ and $p$ are distinct primes. Let $\Pi$ be an $\ell$-modular cuspidal representation of $G(F)$ such that $\Pi$ extends to a representation of $G(F)\rtimes\Gamma$. Then each Tate cohomology group $\widehat{H}^{i}(\Pi)$ is a cuspidal representation of $G^{\sigma}(F)$.
Proof.
Let $Q$ be a proper $F$-parabolic subgroup of $G^{\sigma}$. By [KP23, Lemma 12.5.4], there exists a proper $\sigma$ stable $F$-parabolic subgroup $P$ of $G$ such that $P^{\sigma}=Q$. Let $N$ be the unipotent radical of $P$. Then $N^{\sigma}$ is equal to the unipotent radical of $Q$, to be denoted by $U$. Now, using Theorem 3.3.2 for $\Theta=1$, the trivial character of $N(F)$, we get an injection
$r_{U(F)}(\widehat{H}^{i}(\Pi))\hookrightarrow\widehat{H}^{i}(r_{N(F)}(\Pi)).$
Since $\Pi$ is cuspidal, the Jacquet module $r_{N(F)}(\Pi)$ is trivial, and so is $r_{U(F)}(\widehat{H}^{i}(\Pi))$. Thus, the Tate cohomology group $\widehat{H}^{i}(\Pi)$ is cuspidal. ∎
3.4.
In the paper [TV16, Conjecture 6.1], the authors conjectured that if $\Pi$ is an irreducible $\ell$-modular representation of $G(F)$ such that $\Pi$ extends to a representation of $G(F)\rtimes\Gamma$, then the Tate cohomology groups $\widehat{H}^{i}(\Pi)$, $i\in\{0,1\}$, are finite length representations of $G^{\sigma}(F)$. In our previous work [ND24, Proposition 5.5], this conjecture is proved for $G={\rm GL}_{n}$ where $\sigma$ is induced by a Galois automorphism. In this subsection, we generalise this result for arbitrary connected reductive groups.
3.4.1.
Let $E/F$ be a finite Galois extension with $[E:F]=\ell$, where $\ell$ and $p$ are distinct primes. We fix a generator $\sigma$ of the cyclic group ${\rm Gal}(E/F)$. Let $G$ be a connected reductive group defined over $F$. Fix a minimal $F$-parabolic subgroup $P_{0}$ of $G$. Let $\Omega$ denote the set of proper $F$-parabolic subgroups of $G$ containing $P_{0}$. For $P\in\Omega$, we denote by $U$ the unipotent radical of $P$. Let $\pi$ be a smooth $R$-representation of $G(F)$, where the coefficient field $R$ is either $\overline{\mathbb{Q}}_{\ell}$ or $\overline{\mathbb{F}}_{\ell}$. We have the natural map obtained from Frobenius reciprocity:
$\pi\longrightarrow\prod_{P\in\Omega}i_{P(F)}^{G(F)}\circ r_{U(F)}(\pi),$ (3.3)
and let $\pi_{c}$ be the kernel of the map (3.3). If the Jacquet module $r_{U(F)}(\pi_{c})$ is non-zero for some $P\in\Omega$, then Frobenius reciprocity gives the following non-zero map
$\pi_{c}\longrightarrow\prod_{P\in\Omega}i_{P(F)}^{G(F)}\circ r_{U(F)}(\pi_{c}).$ (3.4)
Since the functor $i_{P(F)}^{G(F)}$ preserves injective maps and $r_{U(F)}$ is an exact functor, we get a contradiction to the fact that $\pi_{c}$ is the kernel of the map (3.3). Thus, $\pi_{c}$ is a cuspidal representation. Using this observation, we now prove the following finiteness result.
Theorem 3.4.1.
Let $E/F$ be a finite Galois extension of prime degree $\ell$, where $\ell\neq p$. Let $G$ be a connected reductive group defined over $F$. Let $\Pi$ be a finite length $\ell$-modular representation of $[G(E)\rtimes{\rm Gal}(E/F)]$. Then, the Tate cohomology groups $\widehat{H}^{i}({\rm Gal}(E/F),\Pi)$ are finite length $\ell$-modular representations of $G(F)$.
Proof.
We use induction on the length of $\Pi$ and the dimension of $G$ to prove the theorem. All Tate cohomology groups are with respect to the action of the Galois group ${\rm Gal}(E/F)$, and we drop it from the notation of Tate cohomology. Suppose $\Pi$ is irreducible. Consider the natural map
$0\longrightarrow\widehat{H}^{i}(\Pi)_{c}\longrightarrow\widehat{H}^{i}(\Pi)\longrightarrow\prod_{P\in\Omega}i_{P(F)}^{G(F)}\circ r_{U(F)}(\widehat{H}^{i}(\Pi))$ (3.5)
For each parabolic $P\in\Omega$ with unipotent radical $U$ and an $F$-rational Levi subgroup $M$, the natural map $\Pi\rightarrow r_{U(E)}(\Pi)$ gives an $M(F)$ equivariant injective map (Theorem 3.3.2)
$r_{U(F)}(\widehat{H}^{i}(\Pi))\hookrightarrow\widehat{H}^{i}(r_{U(E)}(\Pi)).$
Since $r_{U(E)}(\Pi)$ is a finite length representation of $M(E)$ and ${\rm dim}(M)<{\rm dim}(G)$, the Tate cohomology $\widehat{H}^{i}(r_{U(E)}(\Pi))$ has finite length as a representation of $M(F)$ (by induction hypothesis). This implies that $r_{U(F)}(\widehat{H}^{i}(\Pi))$ is a finite length $M(F)$ representation, and therefore $i_{P(F)}^{G(F)}(r_{U(F)}(\widehat{H}^{i}(\Pi)))$ is a finite length $G(F)$ representation as the parabolic induction preserves finite length ([Vig96, Chapter II, Section 5.13]). Thus, the representation
$\prod_{P\in\Omega}i_{P(F)}^{G(F)}\circ r_{U(F)}(\widehat{H}^{i}(\Pi))$
is of finite length. Note that the center $Z(E)$ of $G(E)$ acts on the space of $\Pi$ by a character (as $\Pi$ is irreducible). We deduce that the center $Z(F)$ of $G(F)$ acts on $\widehat{H}^{i}(\Pi)$ by a character. Hence, the cuspidal $G(F)$ representation $\widehat{H}^{i}(\Pi)_{c}$, being a subrepresentation of $\widehat{H}^{i}(\Pi)$, admits a central character. Since $\widehat{H}^{i}(\Pi)$ is finitely generated as a representation of $G(F)$ (Theorem 2.3.3), it follows from [Dat09, Corollary 4.5]) that $\widehat{H}^{i}(\Pi)_{c}$ is also finitely generated. Thus, $\widehat{H}^{i}(\Pi)_{c}$ is a finitely generated $\ell$-modular cuspidal representation of $G(F)$ with a central character and we get that $\widehat{H}^{i}(\Pi)_{c}$ is of finite length ([Vig96, Section 7.4, Chapter I]). We conclude from (3.5) that the $\ell$-modular $G(F)$ representation $\widehat{H}^{i}(\Pi)$ is of finite length.
Now, consider the general case where $\Pi$ is not necessarily an irreducible representation of $\overline{\mathbb{F}}_{\ell}[G(F)]$. Let $\sigma$ be a generator of $\Gamma={\rm Gal}(E/F)$. Let $\Pi_{1}$ be an irreducible sub-representation of $\Pi$ and set $\Pi_{1}^{\prime}$ to be the sub-representation of $\Pi$ spanned by $\Pi_{1}^{\sigma^{i}}$, for $0\leq i<\ell$. Note that $\Pi_{1}^{\prime}$ is $\sigma$-stable, and we consider the short exact sequence of $\overline{\mathbb{F}}_{\ell}[G(E)\rtimes\Gamma]$ modules
$0\longrightarrow\Pi_{1}^{\prime}\longrightarrow\Pi\longrightarrow\Pi/\Pi_{1}^{\prime}\longrightarrow 0$
From the long exact sequence of Tate cohomology, we get the following exact sequence:
$\widehat{H}^{i}(\Pi_{1}^{\prime})\rightarrow\widehat{H}^{i}(\Pi)\rightarrow\widehat{H}^{i}(\Pi/\Pi_{1}^{\prime}).$
If $\Pi_{1}$ is not $\sigma$-stable, then $\widehat{H}^{i}(\Pi_{1}^{\prime})$ is trivial for $i\in\{0,1\}$, and if $\Pi_{1}$ is $\sigma$-stable, then $\widehat{H}^{i}(\Pi_{1}^{\prime})$ is a finite length representation of $G(F)$ from the previous case. Note that the length of $\Pi/\Pi_{1}^{\prime}$ is now strictly less than that of $\Pi$. Then, using induction hypothesis, the Tate cohomology $\widehat{H}^{i}(\Pi/\Pi_{1}^{\prime})$ is a finite length representation of $G(F)$. Thus, we get that $\widehat{H}^{i}(\Pi)$ is a finite length representation of $G(F)$. ∎
As a corollary, we get the following result.
Corollary 3.4.2.
Let $E/F$ be a finite Galois extension of prime degree $\ell$, where $\ell\neq p$. Let $G$ be a connected reductive group defined over $F$. Let $\Pi$ be a finite length integral $\mathcal{K}$ representation of $G(E)\rtimes{\rm Gal}(E/F)$. Then, for any $\Lambda[G(E)\rtimes{\rm Gal}(E/F)]$ stable lattice $\mathcal{L}$ in $\Pi$, the Tate cohomology spaces $\widehat{H}^{i}({\rm Gal}(E/F),\mathcal{L})$ are finite length $\ell$-modular representations of $G(F)$.
Proof.
Recall that $\mathcal{L}$ is a free $\Lambda$-module and that $\mathcal{L}/\ell\mathcal{L}$ is a finite length representation of $G(E)$ (see [Vig96, II.5.11.a] for finiteness). Consider the short exact sequence of $G(E)\rtimes{\rm Gal}(E/F)$ modules
$0\longrightarrow\mathcal{L}\xrightarrow{{\rm mult.}\ell}\mathcal{L}\longrightarrow\mathcal{L}/\ell\mathcal{L}\longrightarrow 0$
By Theorem 3.4.1, the Tate cohomology groups $\widehat{H}^{i}(\mathcal{L}/\ell\mathcal{L})$ are of finite length as representations of $G(F)$. From the long exact sequence of Tate cohomology corresponding to the above short exact sequence, we have
$0\rightarrow\widehat{H}^{0}(\mathcal{L})\rightarrow\widehat{H}^{0}(\mathcal{L}/\ell\mathcal{L})\rightarrow\widehat{H}^{1}(\mathcal{L})\rightarrow 0$
and
$0\rightarrow\widehat{H}^{1}(\mathcal{L})\rightarrow\widehat{H}^{1}(\mathcal{L}/\ell\mathcal{L})\rightarrow\widehat{H}^{0}(\mathcal{L})\rightarrow 0$
Thus, we get that each $\widehat{H}^{i}(\mathcal{L})$ is of finite length. ∎
4. Some genericity and irreducibility results
In this section, $G$ will be a split connected reductive algebraic group defined over $F$. We follow the notations defined in Subsection 3.4.1.
4.0.1.
For this section, we follow [BH02]. Let $B$ be a Borel subgroup of $G$ defined over $F$. Let $E$ be a cyclic Galois extension of $F$ with $[E:F]=\ell$. Let $S$ be a maximal $F$-split torus and let $\Phi$ the set of roots of $S$ in $G$, and let $\Delta$ be the set of simple roots in $\Phi$ for the choice of $B$. Let $\mathfrak{g}$ be the Lie algebra of $G$ and let $\mathfrak{g}_{\alpha}$ be the subspace of $\mathfrak{g}$ for which $S$ acts via $\alpha$. The root subgroup corresponding to $\alpha\in\Delta$ is denoted by $U_{\alpha}$, this is the unique connected subgroup of $G$ whose Lie algebra is $\mathfrak{g}_{\alpha}$. Let $R$ denote $\overline{\mathbb{Q}}_{\ell}$ or $\overline{\mathbb{F}}_{\ell}$. A character $\Theta:U(F)\rightarrow R^{\ast}$ is called non-degenerate whenever $\Theta$ is non-trivial on $U_{\alpha}(F)$ for all $\alpha\in\Delta$. Note that
$U(E)/U(E)_{\rm der}\simeq\prod_{\alpha\in\Delta}V_{\alpha}(E),$
where $V_{\alpha}$ is the one-dimensional $F$-vector group. Thus we choose a non-degenerate character $\Theta_{E}:U(E)\rightarrow R^{\ast}$, which is ${\rm Gal}(E/F)$ stable and the restriction of $\Theta_{E}$ to $U(F)$ is denoted by $\Theta_{F}$ (which is a non-degenerate character of $U(F)$). We set
$\mathfrak{m}_{F}=(U(F),\Theta_{F})\ \text{and}\ \mathfrak{m}_{E}=(U(E),\Theta_{E}).$
The multiplicity one theorem says that
${\rm dim}_{R}\big{(}{\rm Hom}_{G(F)}(V,\,{\rm Ind}_{U(F)}^{G(F)}\Theta_{F})\big{)}\leq 1.$
An irreducible $R$-representation $(\pi,V)$ of $G(F)$ is called $\mathfrak{m}_{F}$-generic if
${\rm dim}_{R}\big{(}{\rm Hom}_{G(F)}(V,\,{\rm Ind}_{U(F)}^{G(F)}\Theta_{F})\big{)}=1.$
An equivalent condition for $(\pi,V)$ to be $\mathfrak{m}_{F}$-generic–which we will use frequently–is that the twisted Jacquet module $V_{U(F),\Theta_{F}}\simeq R$. We also have the obvious notions of $\mathfrak{m}_{E}$-generic representations of $G(E)$. Let $(\pi,V)$ be a smooth generic $R$-representation of $G(F)$. Let $\mathcal{J}_{\pi}$ be a non-zero linear functional in ${\rm Hom}_{U(F)}(V,\Theta_{F})$. Let $\mathbb{W}(\pi,\Theta_{F})\subset{\rm Ind}_{U(F)}^{G(F)}(\Theta_{F})$ be the space of functions $W_{v}:G(F)\rightarrow R$, $v\in V$, where
$W_{v}(g):=\mathcal{J}_{\pi}\big{(}\pi(g)v\big{)},$
for $g\in G(F)$. The map $v\mapsto W_{v}$ induces a $G(F)$ equivariant isomorphism from $(\pi,V)$ onto $\mathbb{W}(\pi,\Theta_{F})$. The space $\mathbb{W}(\pi,\Theta_{F})$ is called the Whittaker model of $(\pi,V)$.
4.0.2.
We now prove that the ${\rm Gal}(E/F)$-Tate cohomology of a generic $\ell$-modular ${\rm Gal}(E/F)$ stable representation of $G(E)$ has a unique generic sub-quotient. We proved this result for the case where $G={\rm GL}_{n}/F$ using Kirillov model. However, we give a simpler proof here. Which will be used in the later sections to prove irreducibility of Tate cohomology groups of certain representations of ${\rm GL}_{n}$ obtained as Base change lifts of cuspidal representations.
Theorem 4.0.1.
Let $E/F$ be a finite Galois extension with $[E:F]=\ell$. Let $G$ be a split reductive algebraic group defined over $F$. Let $(\Pi,V)$ be a ${\rm Gal}(E/F)$ stable $\mathfrak{m}_{E}$-generic $\ell$-modular representation of $G(E)$. Then the Tate cohomology group $\widehat{H}^{0}({\rm Gal}(E/F),\Pi)$ admits a unique $\mathfrak{m}_{F}$-generic sub-quotient.
Proof.
Let $\Gamma$ be the Galois group ${\rm Gal}(E/F)$. Let $\mathcal{J}$ be a non-zero element in ${\rm Hom}_{U(E)}(\Pi,\Theta_{E})$. Let $\mathcal{W}$ be the Whittaker model of $\Pi$. Using Proposition 3.3.2, the Tate cohomology applied to the surjective map $\mathcal{W}\rightarrow\mathcal{W}_{U(E),\Theta_{E}}$ gives the following injection:
$\widehat{H}^{0}(\mathcal{W})_{U(F),\Theta_{F}}\hookrightarrow\widehat{H}^{0}(\mathcal{W}_{U(E),\Theta_{E}}).$
Since ${\rm Gal}(E/F)$ acts trivially on the twisted Jacquet module $\mathcal{W}_{U(E),\Theta_{E}}\simeq\overline{\mathbb{F}}_{\ell}$, the Tate cohomology group $\widehat{H}^{0}(\mathcal{W}_{U(E),\Theta_{E}})$ is a one dimensional vector space over $\overline{\mathbb{F}}_{\ell}$. Now, let us consider the restriction to $G(F)$ map
$\Phi:\mathcal{W}^{\Gamma}\longrightarrow{\rm Ind}_{U(F)}^{G(F)}\Theta_{F}$
$W_{v}\longmapsto{\rm res}_{G(F)}W_{v}.$
It is enough to show that $\Phi$ is non-zero. Suppose $\Phi(W_{v})=0$ for all $W_{v}\in\mathcal{W}^{\Gamma}$. Let $v\in{\rm Ker}(\mathcal{J})\cap\mathcal{W}^{\Gamma}$. Note that $\Pi(g)v\in\operatorname{Ker}(\mathcal{J})$, for $g\in G(F)$ and $g\in U(E)$. This shows that $\Pi(g)v\in{\rm Ker}(\mathcal{J})$, for all $g\in G^{\rm der}(E)$. Also note that ${\rm Ker}(\mathcal{J})$ is stable under the center $Z(E)$, and hence $\langle G(E).v\rangle$ is contained in ${\rm Ker}(\mathcal{J})$. Now, by the irreducibility of $\Pi$, we get a contradiction to the assumption that the map $\Phi$ is non-zero. ∎
4.1. Irreducibility of Tate cohomology for ${\rm GL}_{n}$
Let $G$ be the group ${\rm GL}_{n}$, and we prove the irreducibility of the ${\rm Gal}(E/F)$-Tate cohomology of ${\rm GL}_{n}(E)\rtimes{\rm Gal}(E/F)$ stable lattices in certain integral generic $\mathcal{K}$-representations of ${\rm GL}_{n}(E)$. Assume that $n=\ell m$ for some integer $m\geq 1$. Let $(\Pi,V)$ be the irreducible representation
$\tau\times\tau^{\sigma}\times\cdots\times\tau^{\sigma^{\ell-1}},$ (4.1)
where $\tau$ is an absolutely irreducible cuspidal $\mathcal{K}$-representation of ${\rm GL}_{m}(E)$ such that $\tau$ is not isomorphic to twisted representation $\tau^{\sigma}$. Note that $\Pi\simeq\Pi^{\sigma}$, which induces an action of ${\rm Gal}(E/F)$ on the space $V$ that is compatible with the ${\rm Gal}(E/F)$ action on ${\rm GL}_{n}(E)$. We further assume that $\tau$ is integral, i.e., there exists a ${\rm GL}_{m}(E)$ stable $\Lambda$-lattice in the underlying space of $\tau$. The representation $\Pi$ admits a ${\rm GL}_{n}(E)\rtimes{\rm Gal}(E/F)$ stable $\Lambda$-lattice in $V$. We now prove the following result.
Theorem 4.1.1.
Let $(\Pi,V)$ be an absolutely irreducible, integral, generic $\mathcal{K}$-representation of ${\rm GL}_{n}(E)$ of the form (4.1). Then for any ${\rm GL}_{n}(E)\rtimes{\rm Gal}(E/F)$ stable $\Lambda$-lattice $\mathcal{L}\subseteq V$, the zeroth Tate cohomology group $\widehat{H}^{0}({\rm Gal}(E/F),\mathcal{L})$ is irreducible and cuspidal. Moreover, the first Tate cohomology group $\widehat{H}^{1}({\rm Gal}(E/F),\mathcal{L})$ is trivial.
Proof.
We set $\Gamma={\rm Gal}(E/F)$. To see the cuspidality of the Tate cohomology groups $\widehat{H}^{i}(\Gamma,\mathcal{L})$, consider the Jacquet module $\widehat{H}^{i}(\Gamma,\mathcal{L})_{N(F)}$, where $N$ is the unipotent radical of an $F$-rational proper parabolic subgroup $P$ of ${\rm GL}_{n}$. Let $M$ be an $F$-rational Levi subgroup of $P$. Using Proposition 3.3.2, we have $M(F)$ equivariant injection
$\widehat{H}^{i}(\Gamma,\mathcal{L})_{N(F)}\hookrightarrow\widehat{H}^{i}(\Gamma,\mathcal{L}_{N(E)}).$
We show that the Tate cohomology group $\widehat{H}^{i}(\Gamma,\mathcal{L}_{N(E)})$ is trivial for $i\in\{0,1\}$. Say $W$ is a finite length integral $\ell$-adic $M(E)\rtimes\Gamma$ representation such that the Jordan–Holder factors of $W$, as a representation of $M(E)$, are multiplicity free. Assume that all irreducible $M(E)$ sub-representation of $W$ are not stable under the action of $\Gamma$. The Jacquet module $V_{N(E)}$ is a special case of $W$ and it is integral as $\mathcal{L}_{N(E)}$ is a $M(E)\rtimes\Gamma$ stable $\Lambda$-lattice in $V_{N(E)}$ by [Dat05, Proposition 1.4]. We use induction on the length of $W$ to show that $\widehat{H}^{i}(\Gamma,\mathcal{W})$ is trivial, for all $M(E)\rtimes\Gamma$ stable lattice $\mathcal{W}$ in $W$.
Let $W_{1}$ be an irreducible $M(E)$ sub-representation of $W$ and assume that $W_{1}$ is not fixed by $\Gamma$. The $\Lambda$-module
$\mathcal{M}:=\bigoplus_{i=0}^{\ell-1}\sigma^{i}(\mathcal{W}\cap W_{1})$
is a $M(E)\rtimes\Gamma$ stable lattice in $W_{2}=\bigoplus_{i=0}^{\ell-1}\sigma^{i}(W_{1})$ and it is contained in the lattice $\mathcal{W}$. Since $\mathcal{M}$ is a projective $\Lambda[\Gamma]$-module, the Tate cohomology group $\widehat{H}^{i}(\Gamma,\mathcal{M})$ is trivial for $i\in\{0,1\}$. Now, using the long exact sequence of Tate cohomology corresponding to the following short exact sequence
$0\longrightarrow\mathcal{M}\longrightarrow\mathcal{W}\longrightarrow\mathcal{W}/\mathcal{M}\longrightarrow 0,$
we get that
$\widehat{H}^{i}(\Gamma,\mathcal{W})\simeq\widehat{H}^{i}(\Gamma,\mathcal{W}/\mathcal{M}).$
Since the quotient $\mathcal{W}/\mathcal{M}$ is a $M(E)\rtimes\Gamma$ stable $\Lambda$-lattice in $W/W_{2}$, and the length of $W/W_{2}$ is strictly less than that of $W$, the Tate cohomology group $\widehat{H}^{i}(\Gamma,\mathcal{W}/\mathcal{M})$ is trivial (by induction). Therefore, $\widehat{H}^{i}(\Gamma,\mathcal{W})$ is trivial. Thus, we get that the Tate cohomology $\widehat{H}^{i}(\Gamma,\mathcal{L}_{N(E)})$ is trivial. Hence, the representation $\widehat{H}^{i}(\Gamma,\mathcal{L})$ is cuspidal. The $G(F)$ representation $\widehat{H}^{0}(\Gamma,\mathcal{L})$ admits a unique generic sub–quotient (Theorem 4.0.1). Since $\widehat{H}^{0}(\Gamma,\mathcal{L})$ is also cuspidal, it is irreducible. From the inclusion
$\widehat{H}^{1}(\Gamma,\mathcal{L})_{N(F),\Theta_{F}}\hookrightarrow\widehat{H}^{1}(\Gamma,\mathcal{L}_{N(E),\Theta_{E}}).$
we get that $\widehat{H}^{1}(\Gamma,\mathcal{L}_{N(E),\Theta_{E}})$ is trivial, and so is $\widehat{H}^{1}(\Gamma,\mathcal{L})_{N(F),\Theta_{F}}$. Since $\widehat{H}^{1}(\Gamma,\mathcal{L})$ is also cuspidal, we get that $\widehat{H}^{1}(\Gamma,\mathcal{L})$ is trivial. ∎
Remark 4.1.2.
If $\Pi$ is an absolutely irreducible, integral, generic $\mathcal{K}$-representation of the form (4.1), and the $-\ell$ reduction $r_{\ell}(\Pi)$ is irreducible, then it follows from a similar argument as in Theorem 4.1.1 that $\widehat{H}^{0}({\rm Gal}(E/F),r_{\ell}(\Pi))$ is an irreducible $\ell$-modular cuspidal representation of ${\rm GL}_{n}(F)$.
5. Tate cohomology and Shintani correspondence
In this section, we study ${\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})$-Tate cohomology of Galois invariant lattices in Shintani liftings of cuspidal representations of ${\rm GL}_{n}(\mathbb{F}_{q})$ to ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$. The primary novelty of this section is that it treats the case where the Shintani lifting of a cuspidal representation is not necessarily cuspidal.
5.1. Character formula for finite groups of Lie type
5.1.1.
[…]
5.1.2.
For an $F$-stable maximal torus $T$ and a character $\theta$ of $T^{F}$, Deligne–Lusztig associate a virtual representation $R^{G}_{T,\theta}$. Every irreducible $\ell$-adic representation of $G^{F}$ occurs in a virtual character $R^{G}_{T,\theta}$, for some torus $T$ and $\theta\in\widehat{T^{F}}$. For all unipotent elements $u\in G^{F}$, we denote by $Q_{G,T,F}(u)$, the Green’s function, which is defined as the value $R^{G}_{T,1}(u)$. Let $g\in G^{F}$ be an element with Jordan decomposition $g=su$. Let $C_{G}(s)^{0}$ be the connected component of the centralizer of $s\in G^{F}$. Let $T_{1},T_{2},\dots,T_{k}$ be representatives of the $(C_{G}(s)^{0})^{F}$-conjugacy classes of $F$-stable maximal tori in $C_{G}(s)^{0}$, which are conjugate to $T$ in $G^{F}$. For simplicity, we denote by $H$ the group $C_{G}(s)^{0}$. The Deligne–Lusztig character formula (see [DL76, Theorem 4.2] or [GM20, Lemma 2.2.23, Chapter 2]) gives
$R_{T,\theta}^{G}(su)=\sum_{1\leq i\leq k}Q_{H,T_{i},F}(u)\sum_{w\in W(G,T_{i})^{F}/W(H,T_{i})^{F}}\theta(wsw^{-1}).$ (5.1)
Let $T$ be an $F$-stable maximal torus of $G$ such that $T$ is contained in an $F$-stable parabolic subgroup $P$ of $G$. Let $M$ be an $F$-stable Levi-subgroup of $P$. Let $\theta\in\widehat{T^{F}}$. Then we have
$R_{T,\theta}^{G}={\rm Ind}_{P,M}^{G}(R^{M}_{T,\theta}).$ (5.2)
The notation ${\rm Ind}_{P,M}^{G}$ means the parabolic induction of a virtual character of $M$ to $G$.
5.1.3.
When $G$ is ${\rm GL}_{n}$ and $F$ is the Frobenius endomorphism induced by $x\mapsto x^{q}$, the action of $F$ on $W(G,S_{0})\simeq S_{n}$ is trivial and the $F$-stable maximal tori of $G$ are parametrised by the conjugacy classes of $S_{n}$. Every cuspidal representation of $G^{F}$ is equal to $(-1)^{n-1}R_{T,\theta}^{G}$, where $T$ is the torus corresponding to the conjugacy class of $W(G,S_{0})$ containing $n$-cycles, and $\theta\in\widehat{T^{F}}$ is in general position, i.e., $\theta^{w}\neq\theta$ for all $w\in W(G,T)^{F}$. For an unipotent element $u\in G^{F}$ with associated partition $\mu$ of the integer $n$, and a maximal torus corresponding to a partition $\lambda$ of the integer $n$, the Green’s function $Q_{G,T,F}(u)$ is explicitly computed by J. A. Green, and they are related to Hall polynomials. We will not make use of this relation with Hall polynomials, however, we will use the connection of Green’s polynomials with the representation theory of $S_{n}$ due to T. Springer. In the fundamental work ([Spr76, Theorem 7.8]), T. Springer showed that
$Q_{G,T_{w},F}(u)=\sum_{i\geq 0}a_{i}(w,u)q^{i}$
where the map
$w\mapsto a_{i}(w,u)$
is the character of an irreducible representation of $S_{n}$. These results have their analogues for all finite reductive groups.
5.1.4.
Let $F:{\rm GL}_{n}\rightarrow{\rm GL}_{n}$ be the Frobenius endomorphism of ${\rm GL}_{n}$ induced by the map $x\mapsto x^{q}$ on the entries of the matrices. Let $w\in S_{n}$ be an $n$-cycle, and let $T_{w}$ be an $F$-stable maximal torus of ${\rm GL}_{n}$ corresponding to the element $w$. Let $s\in T_{w}^{F}$ be an element such that the characteristic polynomial of $s$ is $f(X)^{r}$ where $f(X)\in\mathbb{F}_{q}[X]$ is an irreducible polynomial of degree $d$. We have $n=dr$. Let $H$ be the connected group $C_{{\rm GL}_{n}}(s)$. Note that $H(\overline{\mathbb{F}}_{q})$ is isomorphic to ${\rm GL}_{r}(\overline{\mathbb{F}}_{q})^{d}$ and the Frobenius endomorphism $F:H\rightarrow H$ is given by
$(g_{1},g_{2},\dots,g_{d})\longmapsto({\rm Fr}(g_{d}),{\rm Fr}(g_{1}),\dots,{\rm Fr}(g_{d-1})),$
where ${\rm Fr}$ is the Frobenius endomorphism of ${\rm GL}_{r}$ sending a matrix $(g_{ij})\in{\rm GL}_{n}(\overline{\mathbb{F}}_{q})$ to $(g_{ij}^{q})$. Then $H^{F}$ is isomorphic to ${\rm GL}_{r}(\mathbb{F}_{q^{d}})$. Let $D_{r}$ and $B_{r}$ be the diagonal torus and the upper triangular Borel subgroup of ${\rm GL}_{r}$. Note that $D_{r}^{d}$ and $B_{r}^{d}$ give an $F$-stable maximal torus and an $F$-stable Borel subgroup to be denoted by $S_{H}$ and $B_{H}$, respectively. The Frobenius endomorphism $F$ on $H$ induces a map $F^{\ast}$ on $W(H,S_{H})$. If we identify $W(H,S_{H})$ with $S_{r}^{d}$, then
$F^{\ast}(\sigma_{1},\sigma_{2},\dots,\sigma_{d})=(\sigma_{d},\sigma_{1},\dots,\sigma_{d-1}).$
The next lemma is about the $F^{\ell}$-stable maximal tori in $H$ which are conjugate to $T_{w}$ by an element in ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$. Along the way, we prove some standard observations which we could not find a convenient reference.
Lemma 5.1.1.
Let $w$ be an $n$-cycle in $S_{n}$, and let $T_{w}$ be an $F$-stable maximal torus of ${\rm GL}_{n}$ as above. Let $s\in T_{w}^{F}$ be an element with characteristic polynomial $f(X)^{r}$, where $f(X)\in\mathbb{F}_{q}[X]$ is an irreducible polynomial of degree $d$.
  1. (1)
    Let $S_{1}$ and $S_{2}$ be two $F$-stable maximal tori of $H$ such that $gS_{1}g^{-1}=S_{2}$ for some $g\in{\rm GL}_{n}(\mathbb{F}_{q})$. Then there exists an element $h\in H^{F}$ such that $hS_{1}h^{-1}=S_{2}$.
  2. (2)
    Assume that $(\ell,d)=1$. Let $S_{1}$ and $S_{2}$ be two $F^{\ell}$-stable maximal tori of $H$ such that $gS_{1}g^{-1}=S_{2}$ for some $g\in{\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$. Then there exists an element $h\in H^{F^{\ell}}$ such that $hS_{1}h^{-1}=S_{2}$.
  3. (3)
    Let $S$ be an $F$-stable maximal torus of $H$ such that $gSg^{-1}=T_{w}$ for some $g\in{\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$. There exists an element $h\in H^{F^{\ell}}$ such that $hSh^{-1}=T_{w}$.
Proof.
Let $w_{1}$ and $w_{2}$ be two elements of $W(H,S_{H})$ corresponding to $S_{1}$ and $S_{2}$, respectively. We identify the elements $w_{1}$ and $w_{2}$ by the $d$-tuples $(s_{1},\dots,s_{d})$ and $(t_{1},\dots,t_{d})$ respectively, via the isomorphism
$W(H,S_{H})\simeq S_{r}\times\cdots\times S_{r},$
where $s_{i},t_{j}\in S_{r}$ for $1\leq i,j\leq d$. If there exists an element $k\in S_{r}$ such that
$ks_{1}s_{2}\cdots s_{d}k^{-1}=t_{1}t_{2}\cdots t_{d}.$
We define $k_{r}$ by setting
[…]
Let $w=(k_{1},\dots,k_{d-1},k)$ be an element of $W(H,S_{H})$, and we have
$ww_{1}F^{*}(w)^{-1}=w_{2}.$
Thus, $w_{1}$ is $F^{*}$-conjugate to $w_{2}$. This shows that the tuples $(s_{1},\dots,s_{d})$ and $(t_{1},\dots,t_{d})$ are $F^{\ast}$ conjugate in $W(H,S_{H})$ if and only if $s_{1}s_{2}\cdots s_{d}$ and $t_{1}t_{2}\cdots t_{d}$ are conjugate in $S_{r}$. We deduce that any $F^{\ast}$-conjugacy class in $W(H,S_{H})$ has a representative $(s_{1},\dots,s_{d})$ such that $s_{1}=s_{2}=\cdots=s_{d-1}={\rm id}$ and we choose such a representatives for the tuples $(s_{1},s_{2},\dots,s_{d})$ and $(t_{1},t_{2},\dots,t_{d})$. Assume that $s_{d}$ corresponds to the partition $(r_{1},r_{2},\dots,r_{k})$ of $r$. Then $S_{1}$ is an $F$-stable maximal torus of ${\rm GL}_{n}$ corresponding to the partition $(dr_{1},dr_{2},\dots,dr_{k})$. This shows the lemma in this case.
The part (2)(2)( 2 ) of the lemma is similar to part (1)(1)( 1 ). First assume that $(\ell,d)=1$. Note that the Frobenius endomorphism $F$ on $H$ is the composition $\omega\circ{\rm Fr}$, where $\omega$ is the $d$-cycle $(1,2,\dots,d)$. Then we have the identity $F^{\ell}=\omega^{\ell}\circ{\rm Fr}^{\ell}$. Since $\ell$ does not divide $d$, $w^{\ell}$ is also a $d$-cycle. Using a similar argument as in part (1)(1)( 1 ), we get that $S_{1}$ and $S_{2}$ are conjugate in $H^{F^{\ell}}$.
For part (3)(3)( 3 ) of the lemma, using the above discussion for every $(F^{\ast})^{\ell}$-conjugacy class in $W(H,S_{H})$, we can associate $\ell$-partitions $(r_{i1},r_{i2},\dots,r_{ik_{i}})$ of the positive integer $d/\ell$, for $1\leq i\leq\ell$; and the torus associated with this $(F^{\ast})^{\ell}$-conjugacy class of $W(H,S_{H})$ is ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$-conjugate to a torus of ${\rm GL}_{n}$ corresponding to the partition
$(dr_{11}/\ell,dr_{12}/\ell,\dots,dr_{1k_{1}}/\ell,dr_{21}/\ell,dr_{22}/\ell,\dots,dr_{2k_{2}}/\ell,\dots,dr_{\ell 1}/\ell,dr_{\ell 2}/\ell,\dots,dr_{lk_{\ell}}/\ell).$
Now the lemma follows since $k_{i}=1$, for $1\leq i\leq\ell$ and $r_{i1}=r$. This completes the proof. ∎
5.2. Shintani correspondence
Let $\sigma$ be the Frobenius element in ${\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})$. Let $\Pi$ be an $\ell$-adic irreducible representation of ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$ such that $\Pi$ is ${\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})$ equivariant, i.e., $\Pi\simeq\Pi^{\sigma}$. In the article [Shi76, Theorem 1], Shintani proved that for a suitable choice of isomorphism $I_{\sigma}:\Pi\rightarrow\Pi^{\sigma}$, there exists an $\ell$-adic irreducible representation $\pi$ of ${\rm GL}_{n}(\mathbb{F}_{q})$ such that
${\rm Tr}(I_{\sigma}\circ\Pi(g))={\rm Tr}\big{(}\pi({\rm Norm}_{\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}}(g))\big{)},\ \forall g\in{\rm GL}_{n}(\mathbb{F}_{q^{\ell}}),$ (5.3)
where ${\rm Norm}_{\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}}(g)$ is the unique conjugacy class of ${\rm GL}_{n}(\mathbb{F}_{q})$ containing the element
$g\sigma(g)\cdots\sigma^{\ell-1}(g).$
The association $\Pi\mapsto\pi$ gives a bijection from the set of all isomorphism classes of ${\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})$ equivariant irreducible $\ell$-adic representation of ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$ onto the set of isomorphism classes of irreducible $\ell$-adic representations of ${\rm GL}_{n}(\mathbb{F}_{q})$. The inverse of this bijection is called the Shintani base change lifting, denoted by ${\rm bc}_{\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}}$.
5.3. Congruence of characters
In this subsection, we prove certain congruences between the character of a cuspidal representation and the character of its Shintani lifting to ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$, when evaluated at an element of ${\rm GL}_{n}(\mathbb{F}_{q})$. We follow the notations of the preceding subsection 5.2, and let $F$ be the Frobenius endomorphism of ${\rm GL}_{n}/\mathbb{F}_{q}$ induced by the Frobenius automorphism $\overline{\mathbb{F}}_{q}\rightarrow\overline{\mathbb{F}}_{q}$,  $x\mapsto x^{q}$. We begin with the following preliminary lemma. For any representation $\tau$ of ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$, we denote by $\tau^{F}$ the composition of the representation $\tau$ with the automorphism of ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$ induced by $\sigma$. For any representation $\rho$, we denote by $\chi_{\rho}$ its character.
Lemma 5.3.1.
Let $\tau$ be an irreducible $\ell$-adic representation of ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$. Let $\sigma$ be the Frobenius element in ${\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})$, and let $F$ be the Frobenius endomorphism of ${\rm GL}_{n}/\mathbb{F}_{q}$. Let
$\eta=$ $\tau\otimes\tau^{F}\otimes\cdots\otimes\tau^{F^{\ell-1}},$
$\iota(g)=$ $(g,g^{F},\dots,g^{F^{\ell-1}}).$
Then, we have
$\sum_{w\in S_{l}}\chi^{(w)}_{\eta}(\iota(g))\equiv\sum_{i=0}^{\ell-1}\chi_{\tau}(g^{F^{i}})^{\ell}\ \text{mod}\ (\ell),$
for all $g\in{\rm GL}_{m}(\mathbb{F}_{q^{\ell}})$, where $S_{\ell}$ is permutation group of $\ell$ symbols.
Proof.
For any $w\in S_{\ell}$, we have the following character identity
$\chi_{\eta}^{(w)}(g_{0},\dots,g_{\ell-1})=\chi_{\eta}\,(g_{w^{-1}(0)},\dots,g_{w^{-1}(\ell-1)})=\prod_{i=0}^{\ell-1}\chi_{\tau}^{F^{i}}(g_{w^{-1}(i)}),$
for all $g_{i}\in{\rm GL}_{m}(\mathbb{F}_{q^{\ell}})$. Then
$\sum_{w\in S_{\ell}}\chi^{(w)}_{\eta}(\iota(g))=\sum_{w\in S_{\ell}}\Big{(}\prod_{i=0}^{\ell-1}\,\chi_{\tau}(g^{F^{i+w(i)}})\Big{)}.$ (5.4)
Let $X$ denote the set of functions
$\{{\rm id}+w:\mathbb{Z}/\ell\mathbb{Z}\rightarrow\mathbb{Z}/\ell\mathbb{Z}:w\in S_{\ell}\},$
and let $w_{0}$ be the $\ell$-cycle $(0,1,\dots,\ell-1)$. Note that the obvious action of $w_{0}$ on the space of functions $\{f:\mathbb{Z}/\ell\mathbb{Z}\rightarrow\mathbb{Z}/\ell\mathbb{Z}\}$ given by
$(w_{0}f)(x)=f(w_{0}(x))$
preserves the set $X$; this can be seen from the following identity:
$[w_{0}({\rm id}+w)](i)=i+w^{\prime}(i),$
where $w^{\prime}$ is the permutation of $\mathbb{Z}/\ell\mathbb{Z}$ defined by $w^{\prime}=1+ww_{0}$. Now, for $w\in S_{\ell}$, the permutation ${\rm id}+w$ belongs to the fixed point set $X^{w_{0}}$ if and only if there exists an unique integer $k\in\mathbb{Z}/\ell\mathbb{Z}$ such that
[…]
Therefore, from the identity (5.4), we get
$\sum_{w\in S_{\ell}}\chi^{(w)}_{\eta}(\iota(g))=\sum_{i=0}^{\ell-1}\chi_{\tau}(g^{F^{i}})^{\ell}+\sum_{{\rm id}+w\in X\setminus X^{w_{0}}}\Big{(}\prod_{i=0}^{\ell-1}\,\chi_{\tau}(g^{F^{i+w(i)}})\Big{)}.$ (5.5)
Since $\langle w_{0}\rangle$ acts freely on $X\setminus X^{w_{0}}$, there exists a subset $U\subseteq X\setminus X^{w_{0}}$ such that
$X\setminus X^{w_{0}}=\coprod_{r=0}^{\ell-1}w_{0}^{r}U.$
Then we have
$\sum_{{\rm id}+w\in X\setminus X^{w_{0}}}\Big{(}\prod_{i=0}^{\ell-1}\,\chi_{\tau}(g^{F^{({\rm id}+w)(i)}})\Big{)}=\sum_{r=0}^{\ell-1}\,\sum_{{\rm id}+w\in U}\,\prod_{i=0}^{\ell-1}\chi_{\tau}(g^{F^{[w_{0}^{r}({\rm id}+w)](i)}}).$ (5.6)
Recall that $[w_{0}^{r}({\rm id}+w)](i)=(i+r)+w(i+r)$, $0\leq r\leq\ell-1$, and we have
$\prod_{i=0}^{\ell-1}\chi_{\tau}(g^{F^{[w_{0}^{r}({\rm id}+w)](i)}})=\prod_{i=0}^{\ell-1}\chi_{\tau}(g^{F^{i+w(i)}}).$
Using this identity to the equation (5.6), we get
$\sum_{{\rm id}+w\in X\setminus X^{w_{0}}}\Big{(}\prod_{i=0}^{\ell-1}\,\chi_{\tau}(g^{F^{i+w(i)}})\Big{)}=\ell\sum_{{\rm id}+w\in U}\,\prod_{i=0}^{\ell-1}\chi_{\tau}(g^{F^{i+w(i)}}).$
Then it follow from the equation (5.5), we get
$\sum_{w\in S_{\ell}}\chi^{(w)}_{\eta}(\iota(g))\equiv\sum_{i=0}^{\ell-1}\chi_{\tau}(g^{F^{i}})^{\ell}\ \text{mod}\ (\ell).$
Hence the lemma. ∎
5.4.
Let $G\rtimes\Gamma$ be a finite group such that $\Gamma\simeq\mathbb{Z}/\ell\mathbb{Z}$, and we fix a generator $\sigma$ of $\Gamma$. The following lemma, in certain cases, gives the trace of $G^{\sigma}$ action on the Tate cohomology with respect to the action of $\Gamma$. This will be useful when we consider the Tate cohomology of a lattice in Shintani lifts.
Lemma 5.4.1.
Let $G$ be a finite group with an automorphism $\sigma$ of order $\ell$. Let $\Gamma$ be the group generated by $\sigma$. Let $M$ be a $\Lambda[G\rtimes\Gamma]$-module such that $M$ is a free $\Lambda$-module of finite rank. Assume that $\widehat{H}^{1}(\Gamma,M)=0$. Then
${\rm Tr}(g|\widehat{H}^{0}(\Gamma,M))=\overline{{\rm Tr}(g|M)},$
for all $g\in G^{\sigma}$.
Proof.
This lemma was observed by Ronchetti ([Ron16, Theorem 12]). Since, this lemma is used in a crucial way, we give a proof. We may assume that there is a finite unramified extension $K$ of $\mathbb{Q}_{\ell}$ and a $\mathfrak{o}_{K}[G\rtimes\Gamma]$ module $M_{0}$ such that $M_{0}$ is a free $\mathfrak{o}_{K}$ module and $M=M_{0}\otimes_{\mathfrak{o}_{K}}\Lambda$. Assume that $M_{0}=\oplus_{i=1}^{k}N_{i}$, where $N_{i}$ is an indecomposable $\mathbb{Z}_{\ell}[\Gamma]$ submodule of $M$. Since $\widehat{H}^{1}(\Gamma,M)=0$, we get that $N_{i}$ is isomorphic to $\mathbb{Z}_{\ell}$ or $\mathbb{Z}_{\ell}[\Gamma]$. Let $p_{i}:N_{i}\rightarrow M$ and $q_{i}:M\rightarrow N_{i}$ be the embedding and the projection maps, respectively. Note that
${\rm Tr}(g|M)=\sum_{i=1}^{k}{\rm Tr}(q_{i}gp_{i}:N_{i}\rightarrow N_{i}).$
The action of $q_{i}gp_{i}$ on $N_{i}$ has a unique eigenvalue. When $N_{i}=\mathbb{Z}_{\ell}$, the element $q_{i}gp_{i}$ acts by an unit $\xi\in\mathbb{Z}_{\ell}^{\ast}$ whose $-\ell$ reduction $\overline{\xi}$ coincides with the action of $\xi$ on the Tate cohomology space $\widehat{H}^{0}(\Gamma,\mathbb{Z}_{\ell})=\mathbb{Z}/\ell\mathbb{Z}$. Let us consider the case where $N_{i}=\mathbb{Z}_{\ell}[\Gamma]$. Note that $\widehat{H}^{0}(\Gamma,N_{i})=0$. If $\lambda$ is the unique eigenvalue of $q_{i}gp_{i}$ on ${N_{i}}$, then ${\rm Tr}(q_{i}gp_{i})=\ell.\lambda$. Thus,
$\overline{{\rm Tr}(g|_{M})}=0={\rm Tr}(g|_{\widehat{H}^{0}(\Gamma,M)}).$
Hence the lemma. ∎
Proposition 5.4.2.
Let $\Pi$ be an absolutely irreducible generic $\mathcal{K}$ representation of ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$ such that $\Pi\simeq\Pi^{\sigma}$, for all $\sigma\in{\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})$. Let $\pi$ be the $\ell$-adic cuspidal representation of ${\rm GL}_{n}(\mathbb{F}_{q})$ such that $\Pi\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}$ is the Shintani lift of $\pi$. Let $\mathcal{L}$ be a $\Lambda[{\rm GL}_{n}(\mathbb{F}_{q^{\ell}})\rtimes{\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})]$ stable lattice in $\Pi$. Then we have
${\rm Tr}(g|\widehat{H}^{0}({\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}),\mathcal{L}))=\overline{\chi_{\pi}(g)}^{\ell},$
for all $g\in{\rm GL}_{n}(\mathbb{F}_{q})$. (For $x\in\Lambda$, the notation $\overline{x}$ denotes the $-\ell$ reduction of $x$.)
Proof.
Let $\sigma$ be the Frobenius element in ${\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})$ and let $F$ be the corresponding Frobenius endomorphism of ${\rm GL}_{n}$. We fix a ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})\rtimes\langle F\rangle$ stable $\Lambda$-lattice $\mathcal{L}$ in $\Pi$. Since the Tate cohomology group $\widehat{H}^{0}({\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}),\mathcal{L})$ is irreducible and cuspidal (Theorem 4.1.1), it is sufficient to show that
${\rm Tr}(g|{\widehat{H}^{0}({\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}),\mathcal{L})})=\overline{\chi_{\pi}(g)}^{\ell},$ (5.7)
for all $g$ with Jordan decomposition $g=su$, where the characteristic polynomial of $s$ is of the form $f(X)^{r}$ for some irreducible polynomial $f(X)\in\mathbb{F}_{q}[X]$; let $\mathcal{S}$ be the set of all such elements $g\in{\rm GL}_{n}(\mathbb{F}_{q})$. Since the Tate cohomology group $\widehat{H}^{1}({\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}),\mathcal{L})$ is trivial (Theorem 4.1.1), it follows from Lemma 5.4.1 that the equality (5.7) is equivalent to the following identity
$\overline{\chi_{\Pi}(g)}=\overline{\chi_{\pi}(g)}^{\ell},\,g\in\mathcal{S}.$
[…]
$t\mapsto tt^{F}t^{F^{2}}\dots t^{F^{\ell-1}}.$
Let $\eta$ be the composite
[…]
The Shintani lifting of $\pi$ to ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$ is equal to $\Pi=(-1)^{n-\ell}R^{(\ell)}_{T,\eta}$ (see [Kaw87] for a reference). Here $R^{(\ell)}_{T,\eta}$ is the Deligne–Lusztig induction of the character $\eta$ of $T^{F^{\ell}}$ to $G^{F^{\ell}}$.
First we assume that $n=\ell m$ for some integer $m$. Note that $T^{F^{\ell}}$ is identified with $(\mathbb{F}_{q^{n}}^{\ast})^{\ell}$ and $T^{F}$ is identified with $\mathbb{F}_{q^{n}}^{\ast}$ such that ${\rm Nr}$ is given by
$(t_{1},t_{2},\dots,t_{\ell})\mapsto t_{1}t_{2}^{F}\dots t_{\ell}^{F^{\ell-1}},\ t_{i}\in\mathbb{F}_{q^{n}}^{\ast}.$
Let $P$ be an $F^{\ell}$-stable parabolic subgroup containing $T$, and let $M$ be an $F^{\ell}$-stable Levi subgroup of $P$ containing $T$. We can identify the group $M^{F^{\ell}}$ with ${\rm GL}_{n/\ell}(\mathbb{F}_{q^{\ell}})^{\ell}$ and $\Pi=(-1)^{n-\ell}R^{(\ell)}_{T,\eta}$ with the parabolic induction
$\tau\times\tau^{F}\times\cdots\times\tau^{F^{\ell-1}},$
where $\tau$ is the irreducible cuspidal representation of ${\rm GL}_{n/\ell}(\mathbb{F}_{q^{\ell}})$, obtained as the Deligne–Lusztig induction of the character $\theta$ of $\mathbb{F}_{q^{n}}^{\ast}$ (which is considered as $F^{\ell}$-fixed points of a $F^{\ell}$-stable maximal torus of ${\rm GL}_{n/\ell}$ corresponding to an $n/\ell$-cycle).
Let $s\in T^{F}$ be an element with characteristic polynomial $f(X)^{r}$, where $f(X)\in\mathbb{F}_{q}[X]$ is an irreducible polynomial of degree $d$. We have $n=dr$. Let $H$ be the centralizer $C_{{\rm GL}_{n}}(s)$. Let $u\in H^{F}$ be an unipotent element. From the Deligne–Lusztig character formula (5.1) and Lemma 5.1.1, we get that
$\chi_{\Pi}(su)=Q_{H,T,F^{\ell}}(u)\sum_{w\in W(G,T)^{F^{\ell}}/W(H,T)^{F^{\ell}}}\eta(wsw^{-1}).$
We begin with the case where $\ell\mid d$. Let $f=f_{1}f_{2}\cdots f_{\ell}$ be the splitting of $f$ where $f_{i}(X)\in\mathbb{F}_{q^{\ell}}[X]$ is an irreducible polynomial. With the above identification, the element $s\in T^{F}\subset M^{F^{\ell}}$ is of the form
[…]
where $s_{i}$, considered as an element of ${\rm GL}_{n/\ell}(\mathbb{F}_{q^{\ell}})$, has characteristic polynomial $f_{i}(X)^{r}$. Thus, we have
$F(s_{0})=s_{1},F(s_{1})=s_{2},\dots,F(s_{\ell-1})=s_{0}.$
Note that $H^{F^{\ell}}$ can be identified with ${\rm GL}_{r}(\mathbb{F}_{q^{d}})^{\ell}$ and the element $u\in H^{F}\subset H^{F^{\ell}}$ with
$(u_{1},u_{2},\dots,u_{\ell})$
where $u_{i}$ is the unipotent element of ${\rm GL}_{r}(\mathbb{F}_{q^{d}})$ conjugate to $u$. Thus we have
$Q_{H,T,F^{\ell}}(u)=Q_{H,T,F}(u)^{\ell}.$
Let $\gamma$ be the Frobenius element of ${\rm Gal}(\mathbb{F}_{q^{d}}/\mathbb{F}_{q^{\ell}})$, then the set of conjugates
$\{wsw^{-1}:w\in W(G,T)^{F^{\ell}}/W(H,T)^{F^{\ell}}\}$
=$=\{(\gamma^{k_{0}}$ $(s_{\alpha(0)}),\gamma^{k_{1}}(s_{\alpha(1)}),\dots,\gamma^{k_{\ell-1}}(s_{\alpha(\ell-1)}):0\leq k_{i}\leq d/\ell,\,\alpha\in S_{\ell}\}.$
We observe that
$\eta((\gamma^{k_{0}}(s_{\alpha(0)}),\gamma^{k_{1}}(s_{\alpha(1)}),\dots,\gamma^{k_{\ell-1}}(s_{\alpha(\ell-1)})))$
$=\eta((\gamma^{k_{1}}(F(s_{\alpha(1)})),\gamma^{k_{2}}(F(s_{\alpha(2)})),\dots,\gamma^{k_{\ell-1}}(F(s_{\alpha(\ell-1)})),\gamma^{k_{0}}(F(s_{\alpha(0)})))).$
Let $Y$ be the set
$\{(\gamma^{k}(s_{\alpha(0)}),\gamma^{k}(s_{\alpha(1)}),\dots,\gamma^{k}(s_{\alpha(\ell-1)}):0\leq k\leq d/\ell,\,\alpha\in S_{\ell}\}.$
We deduce that
$\sum_{w\in W(G,T)^{F^{\ell}}/W(H,T)^{F^{\ell}}}\eta(wsw^{-1})\equiv\sum_{wsw^{-1}\in\in Y}\eta(wsw^{-1})\ \text{mod}\ (\ell).$
Using Lemma 5.3.1, we get that
$Q_{H,T,F}(u)^{\ell}\sum_{wsw^{-1}\in\in Y}\eta(wsw^{-1})\equiv Q_{H,T,F}(u)^{\ell}\sum_{0\leq i\leq d}\theta(s^{q^{i}})^{\ell}\ \text{mod}\ (\ell).$
This shows the proposition in this case.
We now assume that $(d,\ell)=1$. In this case, we have
$\chi_{\pi}(su)=Q_{H,T,F^{\ell}}(u)\sum_{w\in W(G,T)^{F^{\ell}}/W(H,T)^{F^{\ell}}}\eta(wsw^{-1}).$
Note that
$Y=\{wsw^{-1}:w\in W(G,T)^{F^{\ell}}/W(H,T)^{F^{\ell}}\}=\{(s^{F^{\alpha(0)}},s^{F^{\alpha(1)}},\dots,s^{F^{\alpha(\ell-1)}}):\alpha\in S_{\ell}\}.$
Using Lemma 5.3.1, we get that
$\sum_{y\in Y}\eta(y)=\sum_{0\leq i\leq d}\theta(s^{q^{i}})^{\ell}\ \text{mod}\ (\ell).$
We must now analyse the Green function, for which we first recall the following identifications.
$H^{F}={\rm GL}_{r}(\mathbb{F}_{q^{d}}),\ H^{F^{\ell}}={\rm GL}_{r}(\mathbb{F}_{q^{d\ell}}),T^{F}=\mathbb{F}_{q^{n}}^{\ast},T^{F^{\ell}}=(\mathbb{F}_{q^{n}}^{\ast})^{\ell}.$
[…]
$Q_{H,T,F}(u)=\sum_{i\geq 0}a_{i}(w,u)q^{i},$
and
$Q_{H,T,F^{\ell}}(u)=\sum_{i\geq 0}a_{i}(w^{\ell},u)q^{i\ell},$
where $g\mapsto a_{i}(g,u)$, for $g\in S_{n}$, is a character of a representation of $S_{n}$. Clearly, we have
$a_{i}(w,u)^{\ell}\equiv a_{i}(w^{\ell},u)\ \text{mod}\ (\ell).$
From which we get that
$Q_{H,T,F^{\ell}}(u)\equiv Q_{H,T,F}(u)^{\ell}\ \text{mod}\ (\ell).$
We thus conclude in all cases
$\chi_{\Pi}(su)\equiv\chi_{\pi}(su)^{\ell}\ \text{mod}\ (\ell).$
In the case where $(\ell,n)=1$, the above proposition is proved by Ronchetti ([Ron16, Theorem 13]). For the uniformity of the proof, we treat this case using our method. For $g\in\mathcal{S}$, we have
$R_{T,\eta}^{(\ell)}(g)=Q_{H,T,F^{\ell}}(u)\sum_{w\in W(G,T)^{F^{\ell}}/W(H,T)^{F^{\ell}}}\theta(wsw^{-1})$
and in this case, we have
$H^{F}={\rm GL}_{r}(\mathbb{F}_{q}),\ H^{F^{\ell}}={\rm GL}_{r}(\mathbb{F}_{q^{d\ell}}),\ T^{F}=\mathbb{F}_{q^{\ell}},\ T^{F^{\ell}}=\mathbb{F}_{q^{n\ell}}$
and we have
$Q_{H,T,F^{\ell}}(u)=\sum_{i\geq 0}a_{i}(w,u)q^{i\ell}\equiv(\sum_{i\geq 0}a_{i}(w,u)q^{i})^{\ell}\ \text{mod}\ (\ell).$
Then we have
$\sum_{w\in W(G,T)^{F^{\ell}}/W(H,T)^{F^{\ell}}}\theta(wsw^{-1})=\sum_{i=1}^{d}\theta(s^{q^{i\ell}})\equiv[\sum_{i=1}^{d}\theta(s^{q^{i}})]^{\ell}\ \text{mod}\ (\ell)$
We now deduce our main theorem.
Theorem 5.4.3.
Let $\Pi$ be an absolutely irreducible generic $\mathcal{K}$-representation of ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$. Let $\pi$ be an irreducible $\ell$-adic cuspidal representation of ${\rm GL}_{n}(\mathbb{F}_{q})$ such that $\Pi\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}$ is the Shintani lift of $\pi$. Then for any ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})\rtimes{\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})$ stable $\Lambda$-lattice $\mathcal{L}$ in $\Pi$, the Tate cohomology $\widehat{H}^{0}({\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}),\mathcal{L})$ is isomorphic to the Frobenius twist $r_{\ell}(\pi)^{(\ell)}$.
Proof.
To prove the theorem, we use the fact that two irreducible $\ell$-modular representations of a finite group are isomorphic if and only if they have the same character ([Lam01, Corollary 7.21]). Using Proposition 5.4.2, we get the character identity
${\rm Tr}(g|_{\widehat{H}^{0}({\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}),\mathcal{L})})=\overline{\chi_{\pi}(g)}^{\ell},$
for all $g\in{\rm GL}_{n}(\mathbb{F}_{q})$. Note that the Tate cohomology group $\widehat{H}^{0}({\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}),\mathcal{L})$ and the Frobenius twist $r_{\ell}(\pi)^{(\ell)}$ are both irreducible. Thus, the Tate cohomology $\widehat{H}^{0}({\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}),\mathcal{L})$ is isomorphic to $r_{\ell}(\pi)^{(\ell)}$ as ${\rm GL}_{n}(\mathbb{F}_{q})$ representations. Hence the theorem. ∎
6. Lifting to $p$-adic groups
In this section, we give an application of Theorem 5.4.3 to the depth zero case of irreducible smooth representations of ${\rm GL}_{n}(F)$, where $F$ is a non-Archimedean local field.
6.1. Local base change lifting
For a non-Archimedean local field $K$, let $\mathcal{A}(n,K)$ be the set of isomorphism classes of irreducible $\ell$-adic smooth representations of ${\rm GL}_{n}(K)$. Let $\mathcal{G}(n,K)$ be the set of isomorphism classes of semisimple $n$-dimensional $\ell$-adic Weil–Deligne representations of the Weil group $\mathcal{W}_{K}$. The $\ell$-adic local Langlands correspondence is a bijection
${\rm LLC}_{K}:\mathcal{A}(n,K)\rightarrow\mathcal{G}(n,K),$
which is uniquely characterized by certain identities of local constants that one attaches to the elements of $\mathcal{A}(n,K)$ and $\mathcal{G}(n,K)$ (see [HT01], [Hen00], [Sch13]). Moreover, the set of irreducible $\ell$-adic cuspidal representations of ${\rm GL}_{n}(K)$ are mapped onto the set of $n$-dimensional irreducible $\ell$-adic representations of $\mathcal{W}_{K}$ via ${\rm LLC}_{K}$. Note that the classical local Langlands correspondence is a bijection between the set of isomorphism classes of irreducible smooth complex representations of ${\rm GL}_{n}(K)$ and the set of isomorphism classes of $n$-dimensional complex semisimple Weil–Deligne representations. To get a correspondence over $\overline{\mathbb{Q}}_{\ell}$, one twists the original correspondence by the character $\nu^{(1-n)/2}$, where $\nu$ is the composition of the determinant character of ${\rm GL}_{n}(K)$ and the absolute value of $K$. For details, see [Clo90, Conjecture 4.4, Section 4.2] and [Hen01, Section 7].
Say $F$ be a non-Archimedean local field, and let $E/F$ be a finite Galois extension such that $[E:F]=\ell$. Let $\pi$ be an irreducible $\ell$-adic representation of ${\rm GL}_{n}(F)$. Let $\Pi$ be an irreducible $\ell$-adic representation of ${\rm GL}_{n}(E)$ such that
${\rm res}_{\mathcal{W}_{E}}\big{(}{\rm LLC}_{F}(\pi)\big{)}\simeq{\rm LLC}_{E}(\Pi)$.
The representation $\Pi$ is called the base change lifting of $\pi$, and we write $\Pi={\rm bc}_{E/F}(\pi)$. In this case, $\Pi$ is isomorphic to $\Pi^{\sigma}$ for all $\sigma\in{\rm Gal}(E/F)$.
6.2.
In this part, we discuss the compatibility of the local base change and Shintani lift. Let $K_{0}(F)$ be the compact open subgroup ${\rm GL}_{n}(\mathfrak{o}_{F})$, and let $K_{1}(F)$ be the $p$ subgroup
${\rm Id}_{n}+\varpi_{F}M_{n}(\mathfrak{o}_{F}),$
where $M_{n}(\mathfrak{o}_{F})$ is the ring of $n\times n$ matrices with entries in the ring of integers $\mathfrak{o}_{F}$, and $\varpi_{F}$ is a uniformiser of $F$. The quotient group $K_{0}(F)/K_{1}(F)$ is isomorphic to ${\rm GL}_{n}(\mathbb{F}_{q})$, where $\mathbb{F}_{q}$ is the residue field of $F$. An irreducible smooth representation $(\pi,V)$ of ${\rm GL}_{n}(F)$, where $V$ is a vector space over $\overline{\mathbb{Q}}_{\ell}$ or $\mathcal{K}$ or $\overline{\mathbb{F}}_{\ell}$, is said be of depth-zero if there exists a non-zero vector in $V$ that is fixed under the action of $K_{1}(F)$. Let $\mathcal{A}(n,F)_{0}$ be the set of isomorphism classes of irreducible smooth depth-zero $\ell$-adic representations of ${\rm GL}_{n}(F)$, and let $\mathcal{A}(n,\mathbb{F}_{q})$ be the set of isomorphism classes of irreducible $\ell$-adic representations of ${\rm GL}_{n}(\mathbb{F}_{q})$. In the article [SZ99, Section A.1.3], Schneider–Zink defined a map
$\varphi_{0}^{F}:\mathcal{A}(n,F)_{0}\longrightarrow\mathcal{A}(n,\mathbb{F}_{q}).$
Let $E$ be a finite unramified extension of $F$ with $[E:F]=\ell$. In [SZ08, Appendix A, equation (A.7)], Zink–Silberger proved the following identity
$\varphi_{0}^{E}\circ{\rm bc}_{E/F}={\rm bc}_{\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}}\circ\varphi_{0}^{F},$ (6.1)
which shows the compatibility of local base change with Shintani lifting.
Let $\Gamma$ be the Galois group ${\rm Gal}(E/F)$ and $\sigma$ be a generator of ${\rm Gal}(E/F)$. Let $(\Pi,V)$ be an absolutely irreducible depth-zero integral generic $\mathcal{K}$ representation of ${\rm GL}_{n}(E)$ such that $\Pi\simeq\Pi^{\sigma}$. Let $(\pi,W)$ be the irreducible depth-zero integral $\ell$-adic cuspidal representation of ${\rm GL}_{n}(F)$ such that $\Pi\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}$ is the base change lift of $\pi$. From the definition, we have $\varphi_{0}^{E}(\Pi)=\Pi^{K_{1}(E)}$ and $\varphi_{0}^{F}(\pi)=\pi^{K_{1}(F)}$. Using the identity (6.1), we get that the irreducible ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$ representation $\Pi^{K_{1}(E)}\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}$ is the Shintani lift of the irreducible $\ell$-adic cuspidal representation $\pi^{K_{1}(F)}$ of ${\rm GL}_{n}(\mathbb{F}_{q})$. Using these observations, we now prove the following theorem.
Theorem 6.2.1.
Let $F$ be a non-Archimedean local field with residue characteristic $p$, and let $E/F$ be a finite unramified Galois extension with $[E:F]=\ell$, where $\ell$ and $p$ are distinct odd primes. Let $\Pi$ be a depth-zero, integral, generic, $\mathcal{K}$-representation of ${\rm GL}_{n}(F)$ such that $\Pi$ is absolutely irreducible and $\Pi\simeq\Pi^{\sigma}$. Let $\pi$ be the depth-zero, irreducible, integral, $\ell$-adic cuspidal representation of ${\rm GL}_{n}(F)$ such that $\Pi\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}$ is the base change lifting of $\pi$. Then
  1. (1)
    For any ${\rm GL}_{n}(E)\rtimes{\rm Gal}(E/F)$ stable $\Lambda$-lattice $\mathcal{L}$ of $\Pi$, the Tate cohomology group $\widehat{H}^{0}({\rm Gal}(E/F),\mathcal{L})$ is isomorphic to the Frobenius twist $r_{\ell}(\pi)^{(\ell)}$.
  2. (2)
    If the $-\ell$ reduction $r_{\ell}(\Pi)$ is irreducible, then $\widehat{H}^{0}({\rm Gal}(E/F),r_{\ell}(\Pi))$ is isomorphic to $r_{\ell}(\pi)^{(\ell)}$.
Proof.
Let $\mathcal{V}$ be either the $\Lambda[{\rm GL}_{n}(E)\rtimes\Gamma]$ module $\mathcal{L}$ or the $\overline{\mathbb{F}}_{\ell}[{\rm GL}_{n}(E)\rtimes\Gamma]$ module $r_{\ell}(\Pi)$. Consider the compact open subgroup $K_{1}(E)$ of ${\rm GL}_{n}(E)$. We have the following decomposition
$\mathcal{V}=\mathcal{V}^{K_{1}(E)}\bigoplus\mathcal{V}(K_{1}(E)),$
where $\mathcal{V}(K_{1}(E))=\oplus\mathcal{V}_{\rho}$, where $\rho$ is a non-trivial representation of $K_{1}(E)$ and $\mathcal{V}_{\rho}$ is the $\rho$ isotypic component in $\mathcal{V}$. The space $\mathcal{V}(K_{1}(E))$ is stable under the action of ${\rm Gal}(E/F)$. Using the compatibility of Tate cohomology with finite direct sums, we get
$\widehat{H}^{0}(\mathcal{V})=\widehat{H}^{0}(\mathcal{V}^{K_{1}(E)})\bigoplus\widehat{H}^{0}\big{(}\mathcal{V}(K_{1}(E))\big{)}.$ (6.2)
Note that the ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$ representation $\Pi^{K_{1}(E)}\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}$ is the Shintani lift of the irreducible cuspidal representation $\pi^{K_{1}(F)}$ of ${\rm GL}_{n}(\mathbb{F}_{q})$. Then it follows from Theorem 5.4.3 that $\widehat{H}^{0}(\mathcal{V}^{K_{1}(E)})$ is ${\rm GL}_{n}(\mathbb{F}_{q})$-isomorphic to $r_{\ell}(\pi^{K_{1}(F)})^{(\ell)}$. Since $\widehat{H}^{0}(\mathcal{V})$ is irreducible and cuspidal (Theorem 4.1.1 and Remark 4.1.2), and the space $\widehat{H}^{0}(\mathcal{V})^{K_{1}(F)}$ contains $(r_{\ell}(\pi)^{(\ell)})^{K_{1}(F)}$, we get that $\widehat{H}^{0}(\mathcal{V})$ is isomorphic to $r_{\ell}(\pi)^{(\ell)}$. This completes the proof. ∎
6.3. Tate cohomology of depth-zero generic representations
This part aims to prove Treumann–Venkatesh’s conjecture for $\ell$-adic integral, depth-zero, generic representations of ${\rm GL}_{n}(F)$ without the assumption on the prime $\ell$ made in [ND24], i.e., $\ell$ does not divide $\lvert{\rm GL}_{n-1}(\mathbb{F}_{q})\rvert$ for $n\geq 3$. We first introduce some notations.
6.3.1. Supercuspidal support
Let $\lambda=(n_{1},\dots,n_{r})$ be a partition of $n$. Let $P_{\lambda}$ be the parabolic subgroup of ${\rm GL}_{n}$ consisting of block upper triangular matrices of type $\lambda$. Let $M_{\lambda}$ be the Levi subgroup of $P_{\lambda}$ which is isomorphic to ${\rm GL}_{n_{1}}\times\cdots\times{\rm GL}_{n_{r}}$. For each $1\leq i\leq r$, let $\xi_{i}$ be an irreducible smooth supercuspidal $R$-representation of ${\rm GL}_{n_{i}}(F)$, where $R$ denotes either $\overline{\mathbb{Q}}_{\ell}$ or $\mathcal{K}$ or $\overline{\mathbb{F}}_{\ell}$. Let $\xi$ be the $M_{\lambda}(F)$ representation $\xi_{1}\otimes\cdots\otimes\xi_{r}$. Let $\pi$ be a smooth irreducible $R$-representation of ${\rm GL}_{n}(F)$. There exists a pair $(M_{\lambda}(F),\xi)$ as defined above such that $\pi$ is a sub-quotient of the normalised parabolic induction $i_{P_{\lambda}(F)}^{{\rm GL}_{n}(F)}\xi$. Moreover, the pair $(M_{\lambda}(F),\xi)$ is unique up to ${\rm GL}_{n}(F)$-conjugacy and is called the supercuspidal support of $\pi$. When $R=\overline{\mathbb{Q}}_{\ell}$, this follows from the Frobenius reciprocity (see [BZ77] for instance). When $R=\overline{\mathbb{F}}_{\ell}$, this is a deep result, proved by Vignéras ([Vig98]).
6.3.2.
Let $\pi$ be an $\ell$-adic integral generic representation of ${\rm GL}_{n}(F)$, and let $(M(F),\xi)$ be the supercuspidal support of $\pi$. We assume that $\pi$ is a subrepresentation of $i_{P(F)}^{{\rm GL}_{n}(F)}\xi$, where $P$ is a parabolic subgroup of ${\rm GL}_{n}$ whose Levi component is $M$. Let $\Pi$ be an absolutely irreducible, integral, generic $\mathcal{K}$-representation of ${\rm GL}_{n}(E)$ such that $\Pi\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}$ is the base change lift of $\pi$. Then $\Pi$ is a subrepresentation of $i_{P(E)}^{{\rm GL}_{n}(E)}\eta$ with $\eta=\eta_{1}\otimes\cdots\otimes\eta_{r}$, where $\eta_{i}$ is an absolutely irreducible integral generic $\mathcal{K}$-representation of ${\rm GL}_{n_{i}}(E)$ and $\eta_{i}\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}$ is the base change lift of $\xi_{i}$ for each $1\leq i\leq r$.
6.3.3.
For each $1\leq i\leq r$, let $\mathcal{L}_{i}$ be a $\Lambda[{\rm GL}_{n_{i}}(E)\rtimes\Gamma]$ stable lattice in $\eta_{i}$. Let $\mathcal{L}$ be the module $\mathcal{L}_{1}\otimes\cdots\otimes\mathcal{L}_{r}$, which is a $M(E)\rtimes\Gamma$ stable lattice in $\eta$. By [Vig96, Chapter I, Section 9.3], $i_{P(E)}^{{\rm GL}_{n}(E)}\mathcal{L}$ is a ${\rm GL}_{n}(E)$ stable $\Lambda$-lattice in $i_{P(E)}^{{\rm GL}_{n}(E)}\eta$–which is also stable under the action of $\Gamma$ on $i_{P(E)}^{{\rm GL}_{n}(E)}\eta$, defined by
$(\sigma f)(g)=f(\sigma^{-1}(g))$
for all $\sigma\in\Gamma$, $g\in{\rm GL}_{n}(E)$, and $f\in i_{P(E)}^{{\rm GL}_{n}(E)}\eta$. Let $\mathcal{L}^{\prime}$ be the $\Lambda[{\rm GL}_{n}(E)\rtimes\Gamma]$ stable lattice $\Pi\cap i_{P(E)}^{{\rm GL}_{n}(E)}\mathcal{L}$ in $\Pi$.
Theorem 6.3.1.
Let $\pi$ be a smooth, depth-zero, integral, $\ell$-adic generic representation of ${\rm GL}_{n}(F)$, with $J_{\ell}(\pi)$ the unique generic sub-quotient of the $-\ell$ reduction of $\pi$. Let $\Pi$ be an absolutely irreducible, depth-zero, integral, generic $\mathcal{K}$-representation of ${\rm GL}_{n}(E)$ such that $\Pi\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}$ is the base change lift of $\pi$. Then the Frobenius twist $J_{\ell}(\pi)^{(\ell)}$ is the unique generic sub-quotient of the Tate cohomology group $\widehat{H}^{0}({\rm Gal}(E/F),\mathcal{L}^{\prime})$.
Proof.
Let $(M(F),\xi)$ be the pair such that $\pi$ is a subrepresentation of $i_{P(F)}^{{\rm GL}_{n}(F)}\xi$, where
$M={\rm GL}_{n_{1}}\times\cdots\times{\rm GL}_{n_{r}},\,\,\,\xi=\xi_{1}\otimes\cdots\otimes\xi_{r},$
and each $\xi_{i}$ is an irreducible, integral, $\ell$-adic cuspidal representation of ${\rm GL}_{n_{i}}(F)$. Now, $\Pi\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}$, being the base change lift of $\pi$, is a subrepresentation of $i_{P(E)}^{G_{n}(E)}\eta$ with $\eta=\eta_{1}\otimes\cdots\otimes\eta_{r}$, where each $\eta_{i}$ is an absolutely irreducible, integral, generic $\mathcal{K}$-representation of ${\rm GL}_{n_{i}}(E)$ such that $\eta_{i}\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}$ is the base change lift of $\xi_{i}$. Then we have the following commutative diagram
[…]
where $\varphi$ is a ${\rm GL}_{n}(F)$ equivariant map induced from the injection $\mathcal{L^{\prime}}\hookrightarrow i_{P(E)}^{{\rm GL}_{n}(E)}\mathcal{L}$. Note that the twisted Jacquet modules $\mathcal{L}^{\prime}_{N_{n}(E),\Theta_{E}}$ and $(i_{P(E)}^{{\rm GL}_{n}(E)}\mathcal{L})_{N_{n}(E),\Theta_{E}}$ are free $\Lambda$-modules of rank 111. Also note that the vertical maps $f_{1}$ and $f_{2}$ are non-zero and hence isomorphisms. Therefore, the map $\varphi$ takes the generic subquotients of $\widehat{H}^{0}(\mathcal{L}^{\prime})$ to the generic subquotients of $\widehat{H}^{0}(i_{P(E)}^{{\rm GL}_{n}(E)}\mathcal{L})$. Recall that $\eta_{i}\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}$ is the base change lifting of $\xi_{i}$, and $\mathcal{L}_{i}$ is a ${\rm GL}_{n_{i}}(E)\rtimes{\rm Gal}(E/F)$ stable $\Lambda$-lattice in $\eta_{i}$, for each $1\leq i\leq r$. Then using Theorem 4.1.1, we get that the Tate cohomology group $\widehat{H}^{0}(\mathcal{L}_{i})$ is isomorphic to $r_{\ell}(\xi_{i})^{(\ell)}$, and the equation (2.1) gives the isomorphism
$\widehat{H}^{0}(i_{P(E)}^{{\rm GL}_{n}(E)}\mathcal{L})\simeq i_{P(F)}^{{\rm GL}_{n}(F)}r_{\ell}(\xi)^{(\ell)}.$
Thus, the supercuspidal support of the Frobenius twist $J_{\ell}(\pi)^{(\ell)}$ is same as the supercuspidal support of the generic subquotients of the Tate cohomology $\widehat{H}^{0}(\mathcal{L}^{\prime})$. Hence, $J_{\ell}(\pi)^{(\ell)}$ is a subquotient of $\widehat{H}^{0}(\mathcal{L}^{\prime})$. This completes the proof. ∎
7. The unipotent cuspidal representation of ${\rm Sp}_{4}(\mathbb{F}_{q})$
Finite classical groups can have at most one irreducible unipotent cuspidal representation. Let $G$ be a connected reductive group defined over $\mathbb{F}_{q}$, and let $F:G\rightarrow G$ be the Frobenius endomorphism. Recall that an irreducible representation of a finite reductive group $G^{F}$ is said to be unipotent if it occurs in a Deligne–Lusztig virtual character ±$\pm R_{T,1}$, where $T$ is an $F$-stable torus in $G$. When $G$ is a classical group, if an irreducible unipotent cuspidal representation exists, then it is invariant under automorphisms of the underlying algebraic group. Additionally, we have
  1. (1)
    Unipotent cuspidal representations of finite classical groups are defined over $\mathbb{Q}$ ([Lus02]).
  2. (2)
    The $-\ell$ reduction of an unipotent cuspidal representation is irreducible ([DM18, Theorem 1]).
So, the case of unipotent cuspidal representations for finite classical groups is a natural testing ground for Treumann–Venkatesh conjectures. In this section, we treat the group ${\rm Sp}_{4}(\mathbb{F}_{q})$ which has a unique cuspidal unipotent representation for all $q$. Recall that ${\rm Sp}_{4}(\mathbb{F}_{q})$ is defined as the group of all $g\in{\rm GL}_{4}(\mathbb{F}_{q})$ such that $gJg^{t}=J$ (Note that $g^{t}$ is the transpose of $g$), where
$J=\begin{pmatrix}0&I_{2}\\ -I_{2}&0\end{pmatrix}.$
Let $F$ be the Frobenius endomorphism of ${\rm Sp}_{4}/\mathbb{F}_{q}$. In the famous article [Sri68], B. Srinivasan discovered an irreducible cuspidal representation of the group ${\rm Sp}_{4}(\mathbb{F}_{q})$, which does not have Whittaker model, and this representation is denoted by $\theta_{10}$. The representation $\theta_{10}$ is an unipotent cuspidal representation. In what follows we assume that $\ell\neq 2$.
7.1. Mirabolic restriction of $\theta_{10}$
In [Gol78], Gol’fand proved that $\theta_{10}$, like in the case of irreducible cuspidal representations of ${\rm GL}_{n}(\mathbb{F}_{q})$, has a mirabolic restriction–which is irreducible. To recall these results, we begin by defining some subgroups. Let $P$ be a Siegel parabolic subgroup of ${\rm Sp}_{4}$, whose Levi subgroup $M$ and the unipotent radical $U$ such that
$M(\mathbb{F}_{q})=\bigg{\{}\begin{pmatrix}g&0\\ 0&w\,^{t}g^{-1}w\end{pmatrix}:g\in{\rm GL}_{2}(\mathbb{F}_{q})\bigg{\}},\,\,\,\,U(\mathbb{F}_{q})=\bigg{\{}\begin{pmatrix}I_{2}&X\\ 0&I_{2}\end{pmatrix}:X\in M_{2}(\mathbb{F}_{q}),\,^{t}X=X\bigg{\}},$
where $w=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$ and $I_{2}$ is the $\times 2$ 2 identity matrix. Fix a non-trivial additive character $\psi:\mathbb{F}_{q}\rightarrow\overline{\mathbb{Q}}_{\ell}^{\ast}$, and an element $\tau\in\mathbb{F}_{q}^{\times}\setminus(\mathbb{F}_{q}^{\times})^{2}$. We set
$z=\begin{pmatrix}0&1\\ -\tau&0\end{pmatrix}.$
Consider the character $\psi_{z}:U(\mathbb{F}_{q})\rightarrow\overline{\mathbb{Q}}_{\ell}^{\ast}$, sending
$\begin{pmatrix}I_{2}&X\\ 0&I_{2}\end{pmatrix}\mapsto\psi({\rm Tr}(zX)).$
Let $S(\mathbb{F}_{q})$ be the stabilizer of $\psi_{z}$ in $M(\mathbb{F}_{q})$ and let
$Y(\mathbb{F}_{q})=S(\mathbb{F}_{q})\,U(\mathbb{F}_{q}).$
Then $\psi_{z}$ induces a character $\widetilde{\psi_{z}}:Y(\mathbb{F}_{q})\rightarrow\overline{\mathbb{Q}}_{\ell}^{\ast}$, defined as
$\widetilde{\psi_{z}}(su)={\rm det}(s)\,\psi_{z}(u),$
for $s\in S(\mathbb{F}_{q})$ and $u\in U(\mathbb{F}_{q})$. The induced representation ${\rm Ind}_{Y(\mathbb{F}_{q})}^{P(\mathbb{F}_{q})}(\widetilde{\psi_{z}})$ is irreducible and from [Gol78, Proposition 1] we get that the restriction ${\rm res}_{P(\mathbb{F}_{q})}(\theta_{10})$ is isomorphic to ${\rm Ind}_{Y(\mathbb{F}_{q})}^{P(\mathbb{F}_{q})}(\widetilde{\psi_{z}})$. Moreover, the $-\ell$ reduction of ${\rm Ind}_{Y(\mathbb{F}_{q})}^{P(\mathbb{F}_{q})}(\widetilde{\psi_{z}})$ is irreducible and hence the $-\ell$ reduction $r_{\ell}(\theta_{10})$ is also irreducible.
7.2.
Let $\sigma$ be the Frobenius element in ${\Gamma=\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})$. We denote by $\Pi$ and $\pi$ the unipotent $\ell$-adic cuspidal representation of ${\rm Sp}_{4}(\mathbb{F}_{q^{\ell}})$ and ${\rm Sp}_{4}(\mathbb{F}_{q})$, respectively. By uniqueness of $\Pi$ we have $\Pi\simeq\Pi^{\sigma}$. Let $\eta$ be the additive character of $\mathbb{F}_{q^{\ell}}$ given by the composition $\psi\circ{\rm Tr}_{\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}}$, where ${\rm Tr}_{\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}}$ is the trace function. Since $\ell$ is odd, we have $\tau\in\mathbb{F}_{q^{\ell}}^{\times}\setminus(\mathbb{F}_{q^{\ell}}^{\times})^{2}$. Let $\eta_{z}$ and $\widetilde{\eta}_{z}$ be the corresponding characters of $U(\mathbb{F}_{q^{\ell}})$ and $Y(\mathbb{F}_{q^{\ell}})$, respectively. Then, following the above constructions, we get an irreducible $P(\mathbb{F}_{q^{\ell}})$ representation ${\rm Ind}_{Y(\mathbb{F}_{q^{\ell}})}^{P(\mathbb{F}_{q^{\ell}})}(\widetilde{\eta_{z}})$ such that the restriction ${\rm res}_{P(\mathbb{F}_{q^{\ell}})}(\Pi)$ is isomorphic to ${\rm Ind}_{Y(\mathbb{F}_{q^{l}})}^{P(\mathbb{F}_{q^{\ell}})}(\widetilde{\eta_{z}})$. We use the notations $\rho$ and $\rho_{\ell}$ for the representations ${\rm Ind}_{Y(\mathbb{F}_{q})}^{P(\mathbb{F}_{q})}(\widetilde{\psi_{z}})$ and ${\rm Ind}_{Y(\mathbb{F}_{q^{\ell}})}^{P(\mathbb{F}_{q^{\ell}})}(\widetilde{\eta_{z}})$, respectively.
7.2.1.
Let $X$ and $X_{\ell}$ be the coset spaces $Y(\mathbb{F}_{q})\backslash P(\mathbb{F}_{q})$ and $Y(\mathbb{F}_{q^{\ell}})\backslash P(\mathbb{F}_{q^{\ell}})$, respectively. Since $Y(\mathbb{F}_{q^{\ell}})$ is $\Gamma$ stable subgroup of $P(\mathbb{F}_{q^{\ell}})$, we have the long exact sequence of non-abelian cohomology ([Ser, Appendix, Proposition 1])
$0\rightarrow Y(\mathbb{F}_{q})\rightarrow P(\mathbb{F}_{q})\rightarrow X_{\ell}^{\sigma}\rightarrow H^{1}(\Gamma,Y(\mathbb{F}_{q^{\ell}}))\rightarrow H^{1}(\Gamma,P(\mathbb{F}_{q^{\ell}}))$ (7.1)
Note that $S(\mathbb{F}_{q^{\ell}})$ is a $\Gamma$ stable subgroup of $Y(\mathbb{F}_{q^{\ell}})$ with $S(\mathbb{F}_{q^{\ell}})^{\sigma}=S(\mathbb{F}_{q})$, and we get the following the long exact sequence of non-abelian cohomology
$0\rightarrow S(\mathbb{F}_{q})\rightarrow Y(\mathbb{F}_{q})\rightarrow U(\mathbb{F}_{q})\rightarrow H^{1}(\Gamma,S(\mathbb{F}_{q^{\ell}}))\rightarrow H^{1}(\Gamma,Y(\mathbb{F}_{q^{\ell}}))\rightarrow H^{1}(\Gamma,U(\mathbb{F}_{q^{\ell}})).$
From the above exact sequence, we observe that $H^{1}(\Gamma,S(\mathbb{F}_{q^{\ell}}))$ is isomorphic to $H^{1}(\Gamma,Y(\mathbb{F}_{q^{\ell}}))$. The group $S$ is precisely given by
$\bigg{\{}\begin{pmatrix}g&0\\ 0&w\,^{t}g^{-1}w\end{pmatrix}\in M(\overline{\mathbb{F}}_{q}):gzg^{\ast}=z\bigg{\}},$
where $g^{\ast}$ is the transposition of $g$ with respect to the second diagonal. Note that the determinant map $S\rightarrow\{\pm 1\}$, given by $g\mapsto{\rm det}(g)$ is surjective. The kernel of this determinant map, denoted by $S_{1}$, is the connected component of $S$. Then we have the following long exact sequence
$0\rightarrow H^{0}(\Gamma,S_{1})\rightarrow H^{0}(\Gamma,S)\rightarrow H^{0}(\Gamma,\mathbb{Z}/2\mathbb{Z})\rightarrow H^{1}(\Gamma,S_{1})\rightarrow H^{1}(\Gamma,S)\rightarrow H^{1}(\Gamma,\mathbb{Z}/2\mathbb{Z}).$
Using [DM91, Proposition 4.2.11], we get that $H^{1}(\Gamma,S_{1})$ is trivial. Since $\ell$ is odd, we have $H^{1}(\Gamma,\mathbb{Z}/2\mathbb{Z})=\{0\}$. Then it follows from the above exact sequence that $H^{1}(\Gamma,S)$ is trivial, and so is $H^{1}(\Gamma,Y(\mathbb{F}_{q^{\ell}}))$. Thus, the exact sequence (7.1) gives the equality $X_{\ell}^{\sigma}=X$.
Theorem 7.2.1.
For any ${\rm Sp}_{4}(\mathbb{F}_{q^{\ell}})\rtimes{\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})$ stable $\Lambda$-lattice $\mathcal{L}\subseteq\Pi$, the first Tate cohomology group $\widehat{H}^{1}({\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}),\mathcal{L})$ is trivial.
Proof.
Since ${\rm res}_{P(\mathbb{F}_{q^{\ell}})}(\Pi)$ is isomorphic to $\rho_{\ell}$, it suffices to show that $\widehat{H}^{1}(\mathcal{L})$ is trivial for any $P(\mathbb{F}_{q^{\ell}})\rtimes{\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})$ stable $\Lambda$-lattice $\mathcal{L}$ in $\rho_{\ell}$. Recall that the action of ${\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})$ on $\rho_{\ell}$ is defined by
$(\sigma f)(x)=f(\sigma^{-1}(x)),\ \ \sigma\in{\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})$
for all $x\in P(\mathbb{F}_{q^{\ell}})$ and $f\in\rho_{\ell}$. Let $\mathcal{L}$ be the $P(\mathbb{F}_{q^{\ell}})\rtimes\Gamma$ stable lattice in $\rho_{\ell}$ consisting of $\Lambda$-valued functions in $\rho_{\ell}$. Since $\widetilde{\eta_{z}}(\sigma(y))=\widetilde{\eta_{z}}(y)$ for $y\in Y(\mathbb{F}_{q})$, the Tate cohomology group $\widehat{H}^{1}(\widetilde{\eta}_{z})$ is trivial. Then it follows from the isomorphism (2.1) that
$\widehat{H}^{1}(\mathcal{L})=0.$
Since the $-\ell$ reduction $r_{\ell}(\rho_{\ell})$ is irreducible, any $P(\mathbb{F}_{q^{\ell}})$ stable lattice is homothetic to $\mathcal{L}$. Thus, we get that $\widehat{H}^{1}(\mathcal{L})$ is trivial for any $\Lambda[P(\mathbb{F}_{q^{\ell}})\rtimes\Gamma]$ stable lattice $\mathcal{L}$ in $\rho_{\ell}$. ∎
7.2.2.
Next, we show that the zeroth Tate cohomology of the $-\ell$ reduction of $\Pi$ is irreducible and cuspidal.
Lemma 7.2.2.
Let $\Pi$ be the irreducible unipotent cuspidal representation of ${\rm Sp}_{4}(\mathbb{F}_{q^{\ell}})$. Then the Tate cohomology group $\widehat{H}^{0}(r_{\ell}(\Pi))$ is irreducible and cuspidal.
Proof.
Note that $r_{\ell}(\Pi)$ is irreducible. Further, it follows from the exactness of the Jacquet functor and [Vig96, Chapter I, Section 9.3] that $r_{\ell}(\Pi)$ is also cuspidal. By Corollary 3.3.3, we get that $\widehat{H}^{0}(r_{\ell}(\Pi))$ is cuspidal.
To prove the irreducibility of $\widehat{H}^{0}(r_{\ell}(\Pi))$, we consider the following restriction to $P(\mathbb{F}_{q})$ map
$\big{(}{\rm Ind}_{Y(\mathbb{F}_{q^{\ell}})}^{P(\mathbb{F}_{q^{\ell}})}(r_{\ell}(\widetilde{\eta_{z}}))\big{)}^{\sigma}\longrightarrow{\rm Ind}_{Y(\mathbb{F}_{q})}^{P(\mathbb{F}_{q})}(r_{\ell}(\widetilde{\psi_{z}})^{\ell})$
$f\longmapsto{\rm res}_{P(\mathbb{F}_{q})}(f).$
The above map factors through $\widehat{H}^{0}(r_{\ell}(\rho_{\ell}))$ and induces the following $P(\mathbb{F}_{q})$ isomorphism (see Subsection 2.2)
$\widehat{H}^{0}(r_{\ell}(\rho_{\ell}))\simeq r_{\ell}(\rho)^{(\ell)}.$
Since the restriction ${\rm res}_{P(\mathbb{F}_{q^{\ell}})}(r_{\ell}(\Pi))$ is isomorphic to $r_{\ell}(\rho_{\ell})$, the Tate cohomology $\widehat{H}^{0}(r_{\ell}(\Pi))$, being $P(\mathbb{F}_{q})$ isomorphic to $r_{\ell}(\rho)^{(\ell)}$, is an irreducible representation of $P(\mathbb{F}_{q})$, and hence of ${\rm Sp}_{4}(\mathbb{F}_{q})$. This shows the lemma. ∎
7.3.
Next, we show that the Tate cohomology group $\widehat{H}^{0}(r_{\ell}(\Pi))$ is isomorphic to $r_{\ell}(\pi)^{(\ell)}$. Let $Q$ be the Klingen parabolic subgroup of ${\rm Sp}_{4}$ consisting of matrices $(g_{ij})\in{\rm GL}_{4}$ with $g_{21}=g_{31}=g_{42}=g_{43}=0$. Let $L$ be the Levi subgroup of $Q$ and $H$ be the unipotent radical of $Q$. Then $L(\mathbb{F}_{q})$ is isomorphic to $\mathbb{F}_{q}^{\times}\times{\rm SL}_{2}(\mathbb{F}_{q})$, and $H(\mathbb{F}_{q})$ consists of matrices of the form
$h=\begin{pmatrix}1&x&y&z\\ 0&1&0&y\\ 0&0&1&-x\\ 0&0&0&1\end{pmatrix},\,x,y,z\in\mathbb{F}_{q}.$
7.3.1.
Let $\xi$ denote the irreducible representation of $H(\mathbb{F}_{q})$ on the space ¯$\overline{\mathbb{Q}}_{\ell}[\mathbb{F}_{q}]$ consisting of $\overline{\mathbb{Q}}_{\ell}$-valued functions on $\mathbb{F}_{q}$, defined by
$(\xi(h)\varphi)(t)=\psi(z+2ty+xy)\,\varphi(x+t),$ (7.2)
for $h\in H(\mathbb{F}_{q})$ and $\varphi\in\overline{\mathbb{Q}}_{\ell}[\mathbb{F}_{q}]$. Let $A(\mathbb{F}_{q})$ be the subgroup {±$\{\pm I_{4}\}\,{\rm SL}_{2}(\mathbb{F}_{q})H(\mathbb{F}_{q})$, where $I_{4}$ is the $\times 4$ 4 identity matrix. The $H(\mathbb{F}_{q})$ representation $\xi$ is extended to a representation of $A(\mathbb{F}_{q})$ where the actions of {±$\{\pm I_{4}\}$ and ${\rm SL}_{2}(\mathbb{F}_{q})$ are defined as in [Gol78, equations (1)(1)( 1 ) and (2)(2)( 2 )]. The restriction ${\rm res}_{{\rm SL}_{2}(\mathbb{F}_{q})}(\xi)$ splits into two irreducible representations $\xi^{+}$ and $\xi^{-}$ acting on the spaces of even and odd functions in ¯$\overline{\mathbb{Q}}_{\ell}[\mathbb{F}_{q}]$, respectively. We further extend the representation $\xi^{-}$ trivially to the group $A(\mathbb{F}_{q})$ and set $\widetilde{\xi}=\xi\otimes\xi^{-}$. The induced representation ${\rm Ind}_{A(\mathbb{F}_{q})}^{Q(\mathbb{F}_{q})}(\widetilde{\xi})$ is irreducible, which we denote by $\pi_{1}$. Similarly, we have an irreducible representation $\pi_{1,\tau}={\rm Ind}_{A(\mathbb{F}_{q})}^{Q(\mathbb{F}_{q})}(\widetilde{\xi_{\tau}})$, where $\widetilde{\xi_{\tau}}$ is the representation of $A(\mathbb{F}_{q})$ defined analogously to $\widetilde{\xi}$ by replacing the character $\psi(t)$ by $\psi_{\tau}(t)=\psi(\tau t)$. Then the restriction ${\rm res}_{Q(\mathbb{F}_{q})}(\pi)$ is isomorphic to the direct sum $\pi_{1}\oplus\pi_{1,\tau}$ ([Gol78, Proposition 2]).
7.3.2.
Using similar constructions as in the preceding subsection, we get that the restriction ${\rm res}_{Q(\mathbb{F}_{q^{\ell}})}(\Pi)$ is isomorphic to the direct sum $\Pi_{1}\oplus\Pi_{1,\tau}$, where $\Pi_{1}$ and $\Pi_{1,\alpha}$ are the irreducible induced representations ${\rm Ind}_{A(\mathbb{F}_{q^{\ell}})}^{Q(\mathbb{F}_{q^{\ell}})}(\widetilde{\xi}_{\ell})$ and ${\rm Ind}_{A(\mathbb{F}_{q^{\ell}})}^{Q(\mathbb{F}_{q^{\ell}})}(\widetilde{\xi_{\tau}}_{\ell})$ respectively, and the irreducible representations $\widetilde{\xi}_{\ell}$ (resp. $\widetilde{\xi_{\tau}}_{\ell}$) is defined analogously to $\widetilde{\xi}$ (resp. $\widetilde{\xi_{\tau}}$) by replacing the character $\psi$ by $\eta=\psi\circ{\rm Tr}_{\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}}$. Note that $\Pi_{1}$ and $\Pi_{1,\tau}$ are endowed with an action of $\Gamma$, given by
$(\sigma f)(x)=f(\sigma^{-1}(x)),\,\,\forall x\in Q(\mathbb{F}_{q^{\ell}}),\,f\in\Pi_{1}\,(\text{or}\,\,\Pi_{1,\tau}).$
Let $\chi_{\widetilde{\xi}_{\ell}}$ and $\chi_{\widetilde{\xi_{\tau}}_{\ell}}$ be the characters of $\widetilde{\xi}_{\ell}$ and $\widetilde{\xi_{\tau}}_{\ell}$ respectively. Then using the relation (7.2) and [Gol78, Equations (1)(1)( 1 ) and (2)(2)( 2 )], we get the following identity of characters:
$\overline{\chi_{\widetilde{\xi}_{\ell}}(g)}=\overline{\chi_{\widetilde{\xi}}(g)}^{\ell}\,\,\,\,\,\text{and}\,\,\,\,\overline{\chi_{\widetilde{\xi_{\tau}}_{\ell}}(g)}=\overline{\chi_{\widetilde{\xi_{\tau}}}(g)}^{\ell}$ (7.3)
for all $g\in A(\mathbb{F}_{q})$. We now prove the following theorem.
Theorem 7.3.1.
Let $\ell$ be an odd prime number. Let $\Gamma$ be the Galois group ${\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})$, and let $\sigma$ be the Frobenius element in $\Gamma$. Let $\pi$ and $\Pi$ be the irreducible unipotent $\ell$-adic cuspidal representation of ${\rm Sp}_{4}(\mathbb{F}_{q})$ and ${\rm Sp}_{4}(\mathbb{F}_{q^{\ell}})$, respectively. Then the Tate cohomology group $\widehat{H}^{0}(r_{\ell}(\Pi))$ is isomorphic to the Frobenius twist $r_{\ell}(\pi)^{(\ell)}$.
Proof.
Let $\chi_{\Pi}$ and $\chi_{\pi}$ be the characters of $\Pi$ and $\pi$ respectively. We will show that
$\overline{\chi_{\Pi}(g)}=\overline{\chi_{\pi}(g)}^{\ell},$
for all $g\in{\rm Sp}_{4}(\mathbb{F}_{q})$. Let $g=su=us$ be the Jordan decomposition of $g$. Note that either $g$ is contained in a parabolic subgroup of ${\rm Sp}_{4}(\mathbb{F}_{q})$ or $u$ is the identity element and $s$ is regular. If $s$ is a regular semisimple element, we have $\chi_{\Pi}(s)=1=\chi_{\pi}(s)$. Assume that $g$ belongs to a parabolic subgroup $P(\mathbb{F}_{q})$ or $Q(\mathbb{F}_{q})$. If $g\in P(\mathbb{F}_{q})$, then it follows from the arguments of Lemma 7.2.2 that
$\widehat{H}^{0}(r_{\ell}(\Pi))\simeq r_{\ell}(\pi)^{(\ell)}$
as a representation of $P(\mathbb{F}_{q})$. Thus we have
$\overline{\chi_{\Pi}(g)}=\overline{\chi_{\pi}(g)}^{\ell}$
for all $g\in P(\mathbb{F}_{q})$.
Suppose $g\in Q(\mathbb{F}_{q})$. Note that $\chi_{\Pi}=\chi_{\Pi_{1}}+\chi_{\Pi_{1,\tau}}$ and $\chi_{\pi}=\chi_{\pi_{1}}+\chi_{\pi_{1,\tau}}$. Using character formula for induced representations, we have
$\chi_{\Pi_{1}}(g)=\sum_{y\in Y_{\ell},\,ygy^{-1}\in A(\mathbb{F}_{q^{\ell}})}\chi_{\widetilde{\xi}_{\ell}}(ygy^{-1})$ (7.4)
and
$\chi_{\pi_{1}}(g)=\sum_{y\in Y,\,ygy^{-1}\in A(\mathbb{F}_{q})}\chi_{\widetilde{\xi}}(ygy^{-1}),$ (7.5)
where $Y_{\ell}$ and $Y$ denote the right coset spaces $A(\mathbb{F}_{q^{\ell}})\backslash Q(\mathbb{F}_{q^{\ell}})$ and $A(\mathbb{F}_{q})\backslash Q(\mathbb{F}_{q})$ respectively, with $Y_{\ell}^{\sigma}=Y$. Since $\Gamma$ acts freely on $Y_{\ell}\setminus Y$, there exists a set $\mathcal{U}\subseteq Y_{\ell}\setminus Y$ such that $Y_{\ell}\setminus Y$ is the disjoint union of $\sigma^{i}\mathcal{U}$, $0\leq i\leq\ell-1$. Also note that $\chi_{\widetilde{\xi}_{\ell}}(\sigma(a))=\chi_{\widetilde{\xi}_{\ell}}(a)$ for all $a\in\mathbb{A}(\mathbb{F}_{q^{\ell}})$. Then the identity (7.4) becomes
$\chi_{\Pi_{1}}(g)=\sum_{y\in Y,\,ygy^{-1}\in A(\mathbb{F}_{q})}\chi_{\widetilde{\xi}_{\ell}}(ygy^{-1})\,\,+\,\,\ell\sum_{u\in\mathcal{U},\,ugu^{-1}\in A(\mathbb{F}_{q^{\ell}})}\chi_{\widetilde{\xi}_{\ell}}(ugu^{-1}).$
Using the relations (7.3) and taking $-\ell$ reduction to the above equation, we get
$\overline{\chi_{\Pi_{1}}(g)}=\sum_{y\in Y,\,ygy^{-1}\in A(\mathbb{F}_{q})}\overline{\chi_{\widetilde{\xi}}(ygy^{-1})}^{\ell}.$
Now, taking $-\ell$ reduction to the relation (7.5) and comparing it with the above identity, we get
$\overline{\chi_{\Pi_{1}}(g)}=\overline{\chi_{\pi_{1}}(g)}^{\ell}.$
Similarly, we have
$\overline{\chi_{\Pi_{1,\tau}}(g)}=\overline{\chi_{\pi_{1,\tau}}(g)}^{\ell}.$
Thus, we get the character identity
$\overline{\chi_{\Pi}(g)}=\overline{\chi_{\pi}(g)}^{\ell},$
for all $g\in{\rm Sp}_{4}(\mathbb{F}_{q})$. Using [Lam01, Corollary 7.21], we conclude that the Tate cohomology $\widehat{H}^{0}(r_{\ell}(\Pi))$ is isomorphic to $r_{\ell}(\pi)^{(\ell)}$.
8. Conflict of interest statement
On behalf of all authors, the corresponding author states that there is no conflict of interest.
9. Date statement
No associated data